### v1

# A Joint Learning Framework for Bridging Defect Prediction and Interpretation

Abstract
Over the past fifty years, numerous software defect prediction (SDP) approaches have been proposed. However, the ability to explain why predictors make certain predictions remains limited. Explainable SDP has emerged as a promising solution by using explainable artificial intelligence (XAI) methods to clarify the decision-making processes of predictors. Despite this progress, there is still significant potential to enhance the reliability of existing approaches. To address this limitation, we treat defect prediction and the corresponding interpretation as two distinct but closely related tasks and propose a joint learning framework that allows for the simultaneous training of the predictor and its interpreter. The novelty of our approach lies in two main aspects: 1. We design feedback loops that convey the decision-making logic from the predictor to the interpreter. This ensures a high level of conciseness in decision logic and feature engineering for both the predictor and the interpreter, enabling the interpreter to achieve reliable local and global interpretability. 2. We incorporate the interpretation results as a penalty term in the loss function of the joint-learning framework. This not only improves the accuracy of the predictor but also imposes a stronger constraint on the reliability of the interpreter. We validated our proposed method against several existing explainable SDPs across multiple datasets. The results demonstrate its effectiveness in both interpretation and defect prediction. The source code for the proposed method is available at: https://github.com/BugPredictor/software-defect-prediction.git
Index Terms:
software defect prediction, interpretability, knowledge distillation, joint learning framework
I Introduction
For nearly half a century, software defect prediction (SDP) has been a significant research focus in the field of software engineering [1, 2]. Its primary aim is to identify defective code early in the software development process, helping practitioners prioritize quality assurance tasks, especially when resources are limited or deadlines are tight [49]. Inspired by the success of machine learning (ML) in various disciplines, such as computer vision and machine translation, researchers have explored applying different ML algorithms to improve the accuracy of SDP. However, despite numerous efforts, existing SDP models often fail to provide actionable guidance for practitioners, making practical application challenging [51].
Recent studies have highlighted that the limited applicability of SDP in real-world scenarios is closely linked to the lack of interpretability [3, 4]. Interpretability refers to the ability to provide comprehensible insights into the decision-making logic for humans. Without it, developers struggle to understand the logic behind SDP predictions, leading to distrust and reluctance to adopt these techniques [3, 5]. Explaining prediction results can be particularly challenging for ML-based predictors, especially those using deep learning (DL). The complex structures and large number of parameters in such models obscure straightforward interpretation.
To address this issue, researchers have investigated the feasibility of explainable SDP, incorporating explainable artificial intelligence (XAI) into SDP. Efforts in explainable SDP are generally categorized into two groups [6]: global and local interpretations. Global approaches focus on explaining the overall behavior of a defect predictor, offering insights into how the model makes decisions across all instances [7]. Local approaches provide clear evidence for why a given instance is predicted as buggy or clean. The prediction results provide actionable guidance for practitioners to address key questions, such as why a specific prediction was made and how to mitigate associated risks [57]. As a result, practitioners often prefer to use local interpretation techniques in practice [3]. Nevertheless, local techniques often encounter challenges with reliability under different settings [52].
[The interpretation for the SMOTE-based predictor] Refer to caption [The interpretation for the RUS-based predictor] Refer to caption
(a) (a) The interpretations by different sampling techniques
[The first interpretation] Refer to caption [The second interpretation] Refer to caption
(b) (b) Two interpretations produced by LIME
Figure 1: Interpretations of LIME for a code fragment.
To illustrate the challenges associated with local approaches, we take the interpretation for ”AbstractCommand.java” from the Apache ActiveMQ project as an example. The explanations generated by LIME with different samplers (i.e., SMOTE and RUS) are shown in Fig. 1 LABEL:sub@fig_1:LIME_RUS. The explanations generated by applying LIME twice to the same file are shown in Fig. 1 LABEL:sub@fig1:LIME_two. These figures present the rankings of features that contribute to the predictions. Features displayed in red indicate their contribution to predicting buggy code, while those displayed in green indicate their contribution to the prediction of clean code.
In Fig. 1 LABEL:sub@fig_1:LIME_RUS, we can see that although the SDP model with different samplers predicts the file as buggy, their explanations vary significantly. Among the ten features selected by LIME for the SMOTE-based predictor, only two are also selected for the RUS-based predictor. Fig. 1 LABEL:sub@fig1:LIME_two further demonstrates that the two explanations generated by running the same predictor twice show considerable variation. For instance, in the first explanation, seven features contribute to the prediction of buggy code, whereas in the second explanation, the number increases to eight.
To address this issue, we consider defect prediction and the corresponding interpretation as two distinct but strongly related tasks and propose a joint-learning framework in this paper. Specifically, we design multiple feedback loops between the interpreter and predictor to simultaneously achieve reliable interpretation and high prediction accuracy. The specific contributions are as follows:
  • To improve the reliability of interpretation, we revised the loss function of the knowledge distillation (KD) framework and introduced output fidelity loss and feature fidelity loss. This strategy constructs a knowledge transfer loop between the interpreter and predictor, ensuring a high level of conciseness in decision-making logic and feature engineering power.
  • To improve the accuracy of the predictor, we incorporate the interpretation results as a penalty term within the loss function of the joint-learning framework. This enables the predictor to capture more discriminative features guided by the interpretation results while imposing stronger restrictions on the fidelity of the interpreter to the predictor.
  • Extensive experiments are conducted on 20 datasets with an in-depth analysis. The results indicate that the joint-learning framework improves both the reliability of interpretation and the accuracy of the predictor.
The paper is organized as follows: Section II provides an overview of the existing research on SDP and interpretation methods used in SDP. In Section III, we describe our proposed method in detail, while in Section IV, we provide the details of the experiment. In Section V,we discuss the threats to validity. Finally, in Section VI, we present our conclusion and future works.
II Related Work
This section provides an overview of SDP and explores existing techniques for explainable SDP.
II-A Software defect prediction
The primary goal of SDP is to classify code fragments (e.g., method or class) as either clean or buggy. Constructing a typical SDP model involves three steps. The first step is to transform software metrics, such as cohesion, coupling, complexity, encapsulation, inheritance, and size, into corresponding numerical representations [10]. The second step labels code fragments as buggy or clean based on post-release defects. These defects can be collected from a Bug Tracking System (BTS) by connecting bug reports to their bug-fixing activities. Code fragments associated with these bug-fixing activities are regarded as buggy, while all others are labeled as clean. Finally, the labeled code fragments serve as training data for an ML-based classifier, which is then used to predict whether code fragments in a later version are buggy or clean.
Numerous ML algorithms, including Fuzzy self-organizing maps, K-means, SVM, random forest, logistic regression, etc., have been extensively employed in previous research [15, 16, 17, 18, 19, 20, 21]. However, metrics used in SDP consistently exhibit correlations and redundancies [11], compromising the effectiveness of traditional ML-based predictors [12]. To address this issue, various DL models have been investigated to improve SDP performance. The initial DL-based defect predictor, Deeper, introduced by Yang et al. [22], employed a deep belief network to identify discriminative information in the input data. Experimental results showed that Deeper could identify 32.22% more defects than many traditional ML-based approaches. Inspired by Yang’s work, many researchers have proposed different DL-based solutions. Specifically, Hoang et al. proposed a convolutional neural network (CNN)-based defect prediction model [23], which experimental results indicated outperformed Deeper. Qiao et al. proposed another DL-based model called DPNN to predict the number of defects [24]. Experimental results showed a 14% reduction in mean square error and an 8% increase in the squared correlation coefficient. Li et al. proposed another CNN-based defect predictor, DP-CNN [25], which achieved better performance than other methods. As demonstrated, DL has significantly improved the performance of SDP.
To ensure the generalization, we employ the DP-CNN proposed in [25] as the predictor and evaluate the difference in accuracy before and after integrating it into the joint-learning framework.
II-B Interpretation of defect prediction
In many practical application scenarios, practitioners often ask: What is the underlying decision logic of the predictor? What is the reason for predicting a code fragment as buggy? How do different metrics influence the prediction? These questions highlight an urgent requirement for explainable defect prediction. However, highly accurate SDP models (e.g., Random Forest, DL) are often black-box in nature, making them difficult to interpret.
Efforts to improve the interpretability of predictions can be broadly divided into two main categories [6]: global and local approaches. The global approaches offer a high-level view of how inputs influence the model’s predictions and reveal patterns that are consistent across the model’s entire operation, rather than focusing on individual predictions. Many ML algorithms (e.g., decision trees, logistic regression techniques) and statistical methods (e.g., ANOVA, variable importance) fall into this category[26, 27]. These methods share two common aspects: 1. Simulatability: Every step of the decision-making process is comprehensible. For example, decision trees are simulatable since we can observe the state of every node as well as the feature thresholds used to separate samples. 2. Transparency: The decision-making process can be broken down into interpretable steps. For instance, the interpretability of a decision tree is defined by the computational path from the root node to the leaf nodes. However, global approaches cannot provide detailed interpretations for each code fragment, and their relatively simple structure often fails to guarantee optimal predictive accuracy. Consequently, many practical applications have adopted local methods.
In contrast to global approaches, local methods such as LIME [8] and BreakDown [28], usually interpret the prediction of each code fragment as a list of ordered features. Feedback from practitioners indicates that the interpretations produced by local methods can offer more actionable guidance [52]. However, recent studies have reported that local techniques lack reliability under various conditions [30, 31, 52]. Firstly, the interpretations of the same code can be inconsistent when different data sampling techniques are used. In SDP, sampling is frequently used due to the scarcity of buggy codes. It is an effective way to mitigate the effects of imbalanced data on the ML algorithms [53]. In [52], Jiho Shin et al., investigated the consistency of local methods across five commonly used data sampling techniques. Their analysis revealed that both the features and the ranking of features in the interpretation vary significantly when different data sampling techniques are employed. Secondly, the interpretations of a code fragment do not exhibit consistency when multiple interpretations are executed. The underlying intuition of local methods regards the interpreter $g$ and the predictor $f$ as two independent ML algorithms and leverages $g$ to approximate the behavior of $f$ within the local area of an instance $x$, denoted as $\pi_{x}$. In general, local methods cannot guarantee that the $\pi_{x}$ required for multiple interpretations of the same instance are the same. So different $\pi_{x}$ can lead to distinct interpretations for the same code fragment [31]. Finally, the predictions produced by the interpreter and the predictor for the same code fragment are not consistent. To reduce the difficulty of interpretation, many local methods define the interpreter as a simple-structure ML model. Taking the most widely used local interpretation technique LIME as an example, it sets a linear regression model as the interpreter, and the interpretation of SDP is oversimplified as a linear transformation between input metrics and prediction outcomes. Recently, [31] emphasized that attempting to comprehend an intricate model by employing a simple model might be overly optimistic. The inconsistency between the predictions generated by the interpreter and the predictor indicates that the interpretation fails to capture the underlying decision-making logic of the predictor.
II-C Knowledge distillation
Knowledge distillation (KD) is a typical model compression technique that transfers knowledge from a teacher model (e.g., DL models) to a student model (e.g., a shallow neural network) [32, 33]. KD is regarded as a promising framework for improving the reliability of interpretation due to the faithful knowledge transfer channel between the student and teacher models. In the health care domain, [34] leveraged distilled knowledge to discover disease patterns. Similarly, in the business domain, [35] distilled dynamical pricing knowledge from a complex black-box DL model, optimizing revenue while maintaining interpretability. In this paper, we designate the predictor as the teacher model and the interpreter as the student model, and transfer the decision-making knowledge of predictor to interpreter based on the KD principle. The main difference between our approach and existing methods lies in the revised loss functions of KD framework, which enable us to train the interpreter and predictor in a collaborative manner.
III Proposed Method
Fig. 2 provides an overview of our approach, which consists of four modules: metric selection, predictor and interpreter design, joint learning framework, and local interpretation.
Figure 2: The framework of our approach.
III-A Metric selection
We carefully selected 16 metrics for our method after a thorough investigation of existing literature and metric suites, including the works of [10, 36, 37]. The details of metrics are presented in Table I. The selection of these metrics was guided by two primary criterias, ensuring their relevance and effectiveness in the context of defect prediction: (1) Expressiveness: The metrics needed to be sufficiently expressive, enabling them to capture subtle variations in code quality, structure, and complexity between similar pieces of source code. (2) Proven Effectiveness: Each selected metric has been validated in prior studies [17, 22, 38, 39], demonstrating its utility in software defect prediction tasks. By adhering to these criterias, we constructed a dataset that covered a broad spectrum of code attributes, ranging from complexity measures to process-related characteristics.
TABLE I: Metrics suite
Feature Description
WMC Weighted methods per class
DIT Depth of Inheritance Tree
NOC Number of Children
CBO Coupling between object classes
RFC Response for a Class
LCOM Lack of cohesion in methods
NPM Number of Public Methods
DAM Data Access Metric
MOA Measure of Aggregation
MFA Measure of Functional Abstraction
CAM Cohesion Among Methods of Class
IC Inheritance Coupling
CBM Coupling Between Methods
AMC Average Method Complexity
LOC Lines of Code
CC McCabe’s cyclomatic complexity
III-B Defect prediction and interpretation model
In this section, we provide the detailed information about the predictor and interpreter.
III-B1 Defect predictor
To ensure the generality of our method, we chose a DL-based model, DP-CNN [25] as the predictor. It consists of three main components: the convolution layer, max pooling layer, and classification layer. To further improve accuracy, we added an attention layer to the original model. The overall architecture of DP-CNN is presented in Fig. 3.
Figure 3: Model Architecture of DP-CNN.
The working procedure of DP-CNN can be summarized as follows: First, the input defect metrics are fed into the convolution layer to extract correlations between features. Then feature maps are then passed through the max pooling layer, which reduces the dimensionality while retaining the most significant features. To further enhance the model’s ability to focus on critical patterns, the attention layer dynamically assigns weights to the feature maps. Finally, the processed features are passed through a fully connected neural network and a softmax classifier to generate the prediction result. The correlations between metrics are crucial for defect prediction. Two characteristics of the convolution layer, sparse connectivity and shared weights, make it particularly suitable for this task [25]. Sparse connectivity enables CNNs to have a wider receptive field, capturing non-linear correlations between metrics. Shared weights ensure that each CNN kernel has the same parameters, allowing predictors to capture these correlations regardless of their position in the input metrics. The max pooling layer retains the most significant features by selecting the maximum local element in the convolutional output, enhancing the model’s robustness to changes in the position of metrics in the input. The attention layer dynamically assigns weights to the inputs, providing a flexible way to capture correlations between feature maps. The fully connected neural network is responsible for converting the feature maps generated by previous layers into a logit. The softmax block is positioned at the end of the entire neural network, transforming the logits into the probabilities.
III-B2 Interpretation model
Following the KD principle, we employ the soft decision tree (SDT) as the interpreter [40]. The SDT is a variant of the fuzzy decision tree, sharing a similar structure with the ordinary decision tree but differing in the definition of nodes. In SDT, nodes are defined as neurons with learnable weight $W$ and bias $b$. Weights reflect the importance of each feature in the decision-making process at the node. The output of each node is defined as $P_{i}(x_{i})=\sigma(x_{i}W_{i}+b_{i})$, which determines the probability of transitioning to the right subtree. The $\sigma(\cdot)$ denotes the activation function and $x_{i}$ is the input of the $i$-th node. The output at the $l$-th leaf is defined as:
$Q_{k}^{l}=\frac{{\rm exp}(\phi_{k}^{l})}{\sum_{k}{\rm exp}(\phi_{k}^{l})}$ (1)
where $Q_{k}^{l}$ is the probability at the $l$-th leaf of type $k$ defect, and $\phi_{k}^{l}$ is the learned feature at that leaf.
The reasons we use SDT as an interpreter are as follows: Owing to its neural-like node architecture, SDT exhibits comparable feature engineering capabilities to CNN, enabling it to accurately retain the decision-making knowledge of CNN. Furthermore, in comparison to CNN, the complexity of SDT has been reduced, rendering the decision-making process more comprehensible.
III-C Joint learning framework
The architecture of the joint learning framework is presented in Fig. 4, each feedback loop corresponds to a specific loss function. Training stops when the model reaches the maximum number of iterations or meets the early stopping condition. Given training data set $D=\{\left(x_{i},y_{i}\right)_{i=1}^{N}\}$, $x_{i}\in R^{n}$, $y_{i}\in\{0,1\}$, the predictor $f$ and interpreter $g$, the loss function of joint learning framework is formalized as follows:
Figure 4: The architecture of joint learning framework.
$L=\alpha L_{pred}(f,D)+\beta L_{pred}(g,D)+L_{int}(f,g,D)$ (2)
where $\alpha$ and $\beta$ are hyperparameters.
The original loss function of KD consists of two essential parts: the distillation loss $L_{soft}$ (which is equivalent to $L_{int}(f,g,D)$ in Eq. (2)) and the student loss $L_{hard}$ (which is equivalent to $L_{pred}(g,D)$ in Eq. (2))). $L_{soft}$ quantifies the difference between the teacher model and the student model, which serves as a feedback loop for transferring knowledge from the teacher model to the student model. Minimizing $L_{soft}$ will enforce the student model to adopt similar decision logic to the teacher model. $L_{hard}$ evaluates the difference between the student model’s predictions and the ground truth. Minimizing $L_{hard}$ ensures that the student model retains the appropriate decision logic of the teacher model. In this paper, we introduce two key modifications to the existing loss function: (1) added the output fidelity loss $L_{pred}(f,D)$, and (2) revised the $L_{soft}$.
$L_{pred}(f,D)$ shares a similar definition with $L_{pred}(g,D)$ indicating the consistency of the predictor’s output to the ground truth.
$L_{pred}(f,D)=\sum_{i=1}^{|D|}\left(f(x_{i})-y_{i}\right)^{2}$ (3)
$L_{pred}(g,D)=\sum_{i=1}^{|D|}\left(g(x_{i})-y_{i}\right)^{2}$ (4)
We introduce output fidelity loss guided by two critical considerations: (1) Optimizing the predictor and interpreter simultaneously. Unlike the original KD framework, which keeps the parameters of teacher model (predictor) fixed during the distillation process, minimizing the output fidelity loss will enforce the predictor to capture the discriminative features of the input metrics. (2) Improving the fidelity of the interpreter. Within the KD framework, predictor and interpreter are mutually dependent. Improving the accuracy of the predictor is equivalent to indirectly enforcing the interpreter to provide a more accurate explanation of the predictor’s decision logic. Therefore, introducing the output fidelity loss can boost the performance of the predictor and provide a more robust and insightful knowledge based on the KD principle.
Moreover, we revised the loss function $L_{soft}$ (denoted as $L_{int}(f,g,D)$ in Eq. (2)). According to Eq. (5), the $L_{int}(f,g,D)$ involves two parts, $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$. $L_{of}(f,g,D)$ is equivalent to the distillation loss in the original KD framework. The feature fidelity loss $L_{fef}(f,g,D)$ evaluates the similarity between the feature maps generated by predictor and interpreter. Minimizing $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$ can enforce the interpreter to adopt similar decision logic and feature engineering power of predictor.
$L_{int}(f,g,D)=\lambda L_{of}(f,g,D)+\gamma L_{fef}(f,g,D)$ (5)
where $\lambda$ and $\gamma$ are hyperparameters. $L_{of}(f,g,D)$ shares a similar definition with $L_{soft}(f,g,D)$, which ensures the output of $g$ should be similar to that of $f$.
$L_{of}(f,g,T)=-\sum_{i=1}^{|D|}q_{i}^{T}\log(p_{i}^{T})$ (6)
where $q_{i}^{T}=\frac{{\rm exp}(\nicefrac{{z_{i}}}{{T}})}{\sum_{k}^{N}{\rm exp}(\nicefrac{{z_{k}}}{{T}})}$ and $p_{i}^{T}=\frac{{\rm exp}(\nicefrac{{v_{i}}}{{T}})}{\sum_{k}^{N}{\rm exp}(\nicefrac{{v_{k}}}{{T}})}$ are the classification probability of instance $x_{i}$ generated by $g$ and $f$, $z_{i}$ and $v_{i}$ are the logits of $x_{i}$. $T$ is the temperature hyperparameter [40]. Furthermore, the feature fidelity loss $L_{fef}$ is defined as follows:
$L_{fef}=\sum_{i}^{|D|}\left(d(z_{i})-v_{i}\right)^{2}$ (7)
where function $d(\cdot)$ adjusts $z_{i}$ to the same dimension of $v_{i}$.
To further clarify the proposed joint learning framework, we take the PC2 dataset as an example. The data in PC2 involves software metrics and labels. Software metrics serve as the input for both the predictor $f$ and interpreter $g$. Labels indicate whether a code fragment is buggy or clean. The predictor and interpreter were optimized jointly using the loss function defined in Eq. (2). Through the joint training process, the predictor was continuously refined to improve its classification accuracy. This process enables the predictor to effectively distinguish between buggy and clean code fragments. Simultaneously, the interpreter was trained to generate a decision tree that aligns with the predictor’s decision-making logic. Through this process, the predictor was also influenced by the interpreter, as minimizing $L_{int}(f,g,D)$ required the predictor to refine its feature extraction and decision logic to better align with the explanations provided by the interpreter. In Fig. 5, the decision tree generated by the interpreter reflects the optimized decision logic and highlights key decision paths for defect prediction.
Figure 5: The interpretations of the proposed method on PC2 dataset.
III-D Local interpretation
It is obvious that SDT obtains the global interpretability as decision tree does [41], and interpretation for a dataset can be abstracted as a set of “IF-THEN” rules. In this section, we focus on the interpretability for an instance. Given a metric set $A=\{a_{1},a_{2},...a_{d}\}$, the sensitivity of SDT for an instance $x$ on the $a_{i}$ is defined as follows.
$S(a_{i})=\frac{\Delta g_{i}(x)}{\Delta a_{i}}$ (8)
$\Delta g_{i}(x)=\lvert g(x)-g(x+{\rm One}(a_{i})\Delta a_{i})\rvert$ is the changes of the output of interpreter aroused by the perturbation of the metric $a_{i}$. $\Delta a_{i}$ is the standard variance of $a_{i}$. ${\rm One}(a_{i})$ is the onehot encoder of $i$-the metric.
$\Delta a_{i}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(a_{i}-\bar{a})^{2}}$ (9)
$\bar{a}_{i}=\frac{1}{N}\sum_{j}^{N}a_{j}^{i}$ (10)
where $N$ is the number of training data, and $a_{j}^{i}$ is the $i$-the metric of $x_{j}$. According to Eq. (8), we interpret the correlation between the instance and predictive result as the sensitivity of feature perturbation. Given an instance $x_{i}$, if a metric has a greater sensitivity, that means such a metric has a stronger influence on the prediction result.
Since the $\Delta a_{i}$ is a constant and the parameters of interpreter are fixed after training, the sensitivity $S$ can provide a stable interpretation for any instance. The algorithm for local interpretation is presented below:
Algorithm 1 Local Interpretation
0:  interpreter $g$, instance $x$, metric set $A=\{a_{1},a_{2},\dots,a_{d}\}$, and hyperparameters: $\lambda,\gamma$.
0:  Sensitivity index $SI$
1:  for $a_{i}\in A$ do
2:   Calculate $\text{Calculate }\bar{a}_{i}$ Eq. (10)
3:   Calculate $\text{Calculate }\Delta a_{i}$ Eq. (9)
4:   Calculate $\text{Calculate }\Delta g_{i}(x)$
5:   $S=\cup\ S(a_{i})$ Eq. (8)
6:  end for
7:  $SI\leftarrow$ Sorts $S$ in descending order
8:  return $SI$
The loop structure (lines 2 to 5 in the algorithm) calculates the sensitivity of each metric and generates the sensitivity set $S$. Line 7 rearranges the sensitivity set $S$ in descending order and puts the index of $S$ to $SI$.
IV Experiment
To objectively evaluate the performance of our method, we conducted an experiment to quantify the performance of the proposed method and the baselines on 20 datasets, aiming to answer the following research questions:
RQ1: What reliability-related improvements does the proposed method make in terms of consistency and faithfulness?
As mentioned in Section II, many local methods often yield inconsistent interpretations by adopting different data sampling techniques as well as interpreting the same code fragment multiple times. These two factors greatly affect the reliability of explainable SDP. To evaluate the consistency of interpretation, we adopt the Coincidence Degree (CD) in RQ1 [42]. Given an interpreter $g$, the CD of an instance $x_{i}$ is defined as follows :
$CD(x_{i})=\frac{1}{|M|}\left|\bigcap_{j=1}^{|M|}g_{j}(x_{i})\right|$ (11)
where $\bigcap_{j=1}^{|M|}{g_{j}(x_{i})}$ is the intersection of the multiple interpretations for one instance, $|M|$ represents the number of interpretations or the number of data sampling techniques used. $CD-k\%$ indicates the coincidence of the top $k$ metrics. A greater $CD$ indicates greater consistency.
Furthermore, as discussed in Section II, another issue related to the reliability of interpretation is that predictor and interpreter may generate different predictive results for the same instance. To evaluate the consistency between predictor and interpreter, we introduce Fidelity of Interpretation (FI) [42, 43, 44] here. Given dataset $D$, predictor $f$ and interpreter $g$, the fidelity is defined as:
$FI=\frac{1}{|D|}\sum_{(x_{i},y_{i})\in D}\mathbb{I}\left(f(x_{i}),g(x_{i})\right)$ (12)
$\mathbb{I}(\cdot)$ is an indicator function, which generates 1 when $f(x_{i})=g(x_{i})$, 0 otherwise.
In addition to the FI, we argue that the prediction produced by an interpreter should also align with the ground truth. This alignment indicates how much correct decision knowledge is obtained from the predictive model. In this paper, we use the Accuracy of Interpreter (AI) as another indicator to measure this alignment. Given the dataset $D$ and interpreter $g$, we define the accuracy as:
$AI=\frac{1}{|D|}\sum_{(x_{i},y_{i})\in D}\mathbb{I}\left(g(x_{i}),y_{i}\right)$ (13)
where $\left|D\right|$ denotes the cardinality of $D$, $g\left(x_{i}\right)$ is the prediction made by the interpreter for the instance $x_{i}$.
RQ2: What predictive accuracy-related improvements does the proposed method make?
There are two goals in validating the predictor trained by our framework (denoted as Joint_CNN): (1) The joint learning framework has had a positive impact on the multiple evaluation metrics of the predictor, and (2) its predictive performance is comparable to other notable predictors. For the first goal, we compare the F-measure and AUC of Joint_CNN with the DP-CNN trained not by our framework (denoted as Base_CNN). In addition, we incorporate three additional metrics from [58]: Percent of Perfect Cleans (PPC), Percent of Non-Perfect Cleans (PNPC), and False Omission Rate (FOR). These indices offer valuable insights into the framework’s ability to balance classification accuracy with practical project management benefits, such as budget savings, remaining time estimates, and failure rate predictions. For the second goal, we compare the F-measure and AUC of Joint_CNN with SVM, Random Forest (RF), and Deep Belief Network (DBN) based predictors. Additionally, although SVM, RF, and DBN are commonly used for classification tasks, they are not interpretable prediction models. To better demonstrate the effectiveness and advantages of our model, we also included AutoSpearman [13], a widely recognized interpretable prediction model, in the comparison.
The Percent of Pure Cleanings (PPC) measure is defined in [58] as the percentage of correctly predicted negative cases, which indicates the amount of saved budget in the target project. PPC is calculated as:
$PPC=\frac{\left|TN\right|}{\left|nt\right|}$ (14)
where TN represents the number of true negatives and $\left|nt\right|$ denotes the total number of test instances.
The Percent of Non-Perfect Cleans (PNPC) measure is defined in [58] as the percentage representing the sum of true positives, false positives, and false negatives over the total number of instances tested. This measure is defined as follows:
$PNPC=\frac{\left|TP\right|+\left|FP\right|+\left|FN\right|}{\left|nt\right|}$ (15)
where $\left|TP\right|$ represents the number of true positives, $\left|FP\right|$ denotes the number of false positives, and $\left|FN\right|$ indicates the number of false negatives.
The False Omission Rate (FOR) measure is defined in [58] as the ratio of false negatives over the number of predicted cleans. This is used to estimate the percentage of potential failures that may occur in the target software. FOR is defined as:
$FOR=\frac{\left|FN\right|}{\left|TN\right|+\left|FN\right|}$ (16)
where $\left|FN\right|$ represents the number of false negatives, and $\left|TN\right|+\left|FN\right|$ is the total number of predicted cleans.
The F-measure, also referred to as the F1-score, assesses the balance between precision and recall. It is the harmonic mean of these two values. The F-measure is defined as:
$\color[rgb]{0,0,1}F\text{-measure}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (17)
The Area Under the Curve (AUC) is used to evaluate the performance of binary classification models by measuring the area under its Receiver Operating Characteristic (ROC) curve, which plots the trade-off between true positive rate and false positive rate across different thresholds. The AUC measures the effectiveness of a classification model in distinguishing between defective and non-defective instances. A higher AUC indicates better discrimination, where a value closer to 1 reflects excellent predictive performance.
RQ3: What is the global interpretability of the proposed method?
The objectives of RQ3 are threefold: (1) to demonstrate that our approach obtains global interpretability, (2) to verify its superior computational efficiency, and (3) to demonstrate its reliability. Although local interpretation methods explain the decision logic for a single instance, they fail to reveal the complete structure of predictors [55]. Global interpretable methods can fill in the gaps left by local interpretable methods. For the first objective, we aim to demonstrate our approach is a decision tree-like interpretable method. The substantial computational expense is a primary disadvantage of current global interpretation methods [56]. For the second objective, we want to quantify the computational efficiency of our method and SP-Lime [55] by comparing the training time of them. The rationale for using SP-Lime as the baseline is as follows. First, SP-Lime is a representative example of perturbation-based global interpretable methods and most global explainable methods are based on the perturbation algorithm [55]. Second, SP-Lime is the global version of Lime (one of baselines of RQ1). Selecting it as baseline is helpful to maintain the consistency of the experiment.
RQ4: What is the effect of different features on the performance and interpretability of our approach?
To assess the impact of individual features on the model’s performance and interpretability, we carried out an ablation experiment. Specifically, we took Joint_CNN as the predictor and eliminated metrics individually. The impacts on the predictor was evaluated by observing changes in the F-measure, AUC, and MCC, and compared these results with the model’s interpretability. MCC is an acronym for Matthews Correlation Coefficient that is used to measure the quality of binary classifications, particularly in imbalanced datasets. The MCC is defined as follows:
$\color[rgb]{0,0,1}MCC=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (18)
where $TP$, $TN$, $FP$ and $FN$ represent the number of true positives, true negatives, false positives, and false negatives, respectively. The MCC value ranges from -1 to 1, where -1 indicates complete disagreement between predictions and actual outcomes, 1 represents perfect agreement, and 0 implies performance equivalent to random guessing.
IV-A Dataset
The 20 datasets used in the experiment were collected from the AEEEM 111 https://zenodo.org/record/3362613#.YmuGoNpByUk, NASA  [^2] http://promise.site.uottawa.ca/SERepository/datasets-page.html and PROMISE  [^3] http://promise.site.uottawa.ca/SERepository repositories. The effectiveness of these datasets has been demonstrated by numerous studies [45, 46, 47, 48]. Further details about each dataset are provided in Table II. We observe that these datasets exhibit significant class imbalance, with defective instances accounting for only 0.41% of the PC2 dataset.
TABLE II: Details of the datasets used in the experiment
Group Dataset Instances Defects %Defects
AEEEM JDT 997 206 20.66%
LC 691 64 9.26%
ML 1862 245 13.16%
PDE 1492 209 14.01%
NASA kc1 2109 326 15.46%
mc1 9466 68 0.72%
jm1 7782 1672 21.49%
PC1 549 77 14.03%
PC2 5589 23 0.41%
PC3 1563 160 10.24%
PC4 1458 178 12.21%
PROMISE ant-1.7 745 166 22.28%
camel-1.6 965 188 19.48%
ivy-1.2 352 40 11.36%
jedit-4.1 312 79 25.32%
log4j-1.0 135 34 25.19%
lucene-2.4 340 203 59.71%
poi-3.0 442 281 63.57%
synapse1.2 256 86 33.59%
xerces-1.3 453 69 15.23%
IV-B Setting
In the experiment, some parameters were set as follows: the initial learning rate $lr$ = 1e-06 and the batch size $batch\_size$ is set to 16. The depth of SDT $tree\_depth$ is set to 4. The penalty strength of SDT $penalty\_strength$ is set to 1e+1 and the penalty decay rate $penalty\_decay$ is set to 0.25. The hyperparameters $\gamma$, $\lambda$, $\alpha$ and $\beta$ are set to 0.8, 0.5, 1.4 and 0.6 respectively. The temperature hyperparameter T is set to 100. The window size for the Exponential Moving Average (EMA) $ema\_win\_size$ is set to 1000. We applied the same set of parameters across all datasets to ensure a fair comparison. The dataset was split into 7:2:1 for training, testing, and validation. The experiments are implemented by TensorFlow 2.3.0 with the NVIDIA GeForce RTX 3090 GPU.
IV-B1 Baseline
We utilize LIME [8] and BreakDown [28] as baselines for RQ1 and RQ3. We selected these techniques due to the following reasons: (1) These two methods are extensively employed by the SDP research community; (2) LIME, BreakDown, and our approach belong to feature-based interpretation techniques that strive to quantify the correlations between input features and predictive outputs; (3) The theoretical foundations of LIME and BreakDown differ substantially, thereby enriching the diversity of comparisons. LIME approximates the behavior of predictors in a local context, whereas BreakDown calculates the cumulative contribution of each feature to the prediction result.
IV-B2 Data preprocessing
We leverage normalization and different data sampling techniques to process the imbalanced data. Specifically, normalization is defined as follows:
$\widetilde{x}=\frac{x-\min{\left(D\right)}}{\max{\left(D\right)}-\min{\left(D\right)}}$ (19)
where $x$ is the sample of a dataset, and ${\rm min}(\cdot),{\rm max}(\cdot)$ are the minimum and maximum operations applied column-wise across the dataset. The features in the dataset exhibit significant differences in value range. Without normalization, features with larger numerical ranges might unfairly influence the model. Normalization ensures that all features are on the same scale, so the model treats them equally in terms of their impact. Moreover, normalization helps the algorithm converge faster and more smoothly. As shown in the Table III, normalizing the data significantly improves the model’s performance compared to using the raw features. We adopt two sampling techniques to the dataset: SMOTE (Synthetic Minority Oversampling Technique) and RUS (Random Undersampling). The technical details of SMOTE and RUS can be found in [49] and [54]. These techniques address class imbalance in defect prediction tasks but may also affect model interpretability. SMOTE balances the dataset by generating synthetic samples, which may subtly alter the distribution of dataset and distort the contribution of feature to the prediction results. In contrast, RUS reduces the dataset size by randomly removing majority-class samples, potentially simplifying the decision process but at the risk of discarding important information, which may affect the stability of the generated explanations.
TABLE III: Average Performance of the Models with and without Data Preprocessing Across All Datasets
F-measure AUC
Raw Data 0.5822 0.6041
Preprocessed Data 0.6704 0.7109
IV-B3 Statistical significance test
To verify the significance of the improvements achieved by our approach, we introduced $Cohen^{\prime}s\ d$ in this section. $Cohen^{\prime}s\ d$ is used to measure the difference between two groups, which can take any value between 0 and infinity. When $Cohen^{\prime}s\ d$ takes values in the ranges $[0,0.2),[0.2,0.8),and\ [0.8,+\infty)$, it indicates that the differences between the two groups are small, medium, and large 444 https://en.wikipedia.org/wiki/Effect_size. The definition of $Cohen^{\prime}s\ d$ is as follows:
$Cohen^{\prime}s\ d=\frac{M_{1}-M_{2}}{\sqrt{(SD_{1}^{2}+SD_{2}^{2})/2}}$ (20)
where $M_{1}$ and $M_{2}$ represent the means of the two group data, and $SD_{1}$ and $SD_{2}$ represent their standard deviations.
TABLE IV: The AI and FI of Joint_SDT and LIME
$AI$ $FI$
LIME Joint_SDT LIME Joint_SDT
JDT 0.8200 0.8200 0.8100 0.9300
LC 0.8143 0.8429 0.8000 0.8714
ML 0.6684 0.9037 0.6791 0.9786
PDE 0.7200 0.7533 0.6533 0.7867
jm1 0.6239 0.8010 0.8793 0.9910
kc1 0.6540 0.8531 0.5355 0.9431
mc1 0.8754 0.9916 0.9155 0.9979
PC1 0.8727 0.8727 0.7091 0.9273
PC2 0.9821 0.9982 0.9714 0.9803
PC3 0.7834 0.9172 0.8471 0.9682
PC4 0.8151 0.9178 0.9041 0.9932
ant-1.7 0.6667 0.7467 0.6667 0.9200
camel1.6 0.7010 0.5361 0.7010 0.8144
ivy-1.2 0.8056 0.8333 0.8056 0.9167
jedit-4.1 0.5312 0.7500 0.5312 0.5938
log4j-1.0 0.7857 0.6429 0.7857 0.5714
lucene-2.4 0.5000 0.7059 0.5000 0.7059
poi-3.0 0.4222 0.7778 0.4222 0.8222
synapse-1.2 0.5769 0.6538 0.5769 0.5769
xerces-1.3 0.6522 0.7174 0.6522 0.8478
Average 0.7135 0.8018 0.7173 0.8568
Improvement 12.38%↑ - 19.45%↑ -
$Cohen^{\prime}s\ d$ 0.6792 - 0.9451 -
TABLE V: Average CDs of LIME, BreakDown and proposed method
LIME BreakDown Joint_SDT
Average 65% 74.5% 100%
Improvement 53.85%↑ 34.23%↑ -
IV-C Experimental results
The experimental results are presented as follows.
IV-C1 The results of RQ1
To evaluate the consistency of the interpreter trained by our framework (denoted as Joint_SDT), LIME and BreakDown, we employed each of them to interpret an instance ten times, and calculate the average CD for each method. As shown in Table V, the average CD-10% of LIME, BreakDown and Joint_SDT are 65%, 74.5%, and 100%. This result suggests that LIME exhibits poor consistency. To further investigate the consistency of LIME, we visualized its outputs in Fig. 6 and calculated the coefficient of variation (CV) for the weights assigned to different metrics across the ten interpretations. The results show that the CV is 42.18%. The high CV suggests that LIME is unable to consistently assign similar weights to the same metric across multiple explanations. We believe this inconsistency is partly due to the sampling technique and the hyperparameters used by LIME.
(a) The first interpretation
(b) The second interpretation
(c) The third interpretation
(d) The fourth interpretation
(e) The fifth interpretation
(f) The sixth interpretation
(g) The seventh interpretation
(h) The eighth interpretation
(i) The ninth interpretation
(j) The tenth interpretation
Figure 6: Multiple interpretations of LIME for a certain instance in the PC2 dataset.
Furthermore, we conducted a comparison between the interpretations generated by LIME and Joint_SDT in terms of AI and FI. Since BreakDown does not provide an explicit surrogate function like LIME [8], its AI and FI of BreakDown cannot be calculated. As shown in Table IV, across all projects, Joint_SDT achieves an average improvement of 12.38% in AI and 19.45% in FI compared to LIME. The $Cohen^{\prime}s\ d$ results in the table further confirm that Joint_SDT outperforms LIME, with a large effect size in AI and a moderate effect size in FI.
IV-C2 The results of RQ2
TABLE VI: Prediction Performance of Joint_CNN and Base_CNN in Terms of F-measure and AUC
F-measure AUC
Base_CNN Joint_CNN Base_CNN Joint_CNN
JDT 0.6802 0.8544 0.7297 0.8360
LC 0.5686 0.8285 0.6923 0.6862
ML 0.5686 0.7123 0.6897 0.6418
PDE 0.7475 0.7111 0.7039 0.7032
jm1 0.5755 0.8621 0.6240 0.6280
kc1 0.6443 0.7303 0.7545 0.8087
mc1 0.7352 0.8621 0.8082 0.6280
PC1 0.7475 0.8621 0.7039 0.6280
PC2 0.4996 0.8621 0.5000 0.6280
PC3 0.5588 0.7957 0.7634 0.8324
PC4 0.7021 0.7472 0.8169 0.7843
ant-1.7 0.7991 0.7907 0.8276 0.8022
camel-1.6 0.6932 0.7337 0.5992 0.6687
ivy-1.2 0.8273 0.8347 0.6016 0.8594
jedit-4.1 0.5862 0.6946 0.5429 0.8125
log4j-1.0 0.7755 0.6520 0.8500 0.8500
lucene-2.4 0.5623 0.6947 0.5482 0.7643
poi-3.0 0.7156 0.6951 0.7263 0.7414
synapse1.2 0.6616 0.6486 0.8235 0.7516
xerces-1.3 0.7921 0.8407 0.8498 0.7875
Average 0.6720 0.7706 0.7078 0.7421
Improvement 14.67%↑ - 4.85%↑ -
$Cohen^{\prime}s\ d$ 1.1321 - 0.3553 -
TABLE VII: Average Performances of Joint_CNN and Base_CNN in Terms of PPC, PNPC and FOR
Group PPC PNPC FOR
Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN
AEEEM 0.6570 0.7394 0.3430 0.2606 0.2597 0.1662
NASA 0.6811 0.7708 0.3189 0.2292 0.0503 0.0256
PROMISE 0.4974 0.5167 0.5026 0.4833 0.1867 0.1398
Average 0.6118 0.6756 0.3882 0.3244 0.1656 0.1105
Improvement 10.43% - 19.67$\downarrow$ - 49.86$\downarrow$ -
According to Table VI, we observed that Joint_CNN outperforms Base_CNN in terms of both F-measure and AUC. Across all datasets, Joint_CNN achieves an average improvement of 14.67% in F-measure and 4.85% in AUC compared to Base_CNN. Table VII presents the average performance of Base_CNN and Joint_CNN on three groups: PROMISE, NASA, and AEEEM, evaluated using the metrics PPC, PNPC, and FOR. The results in Table VII show that, compared to Base_CNN, Joint_CNN achieves improvements of 10.43% in PPC, 19.67% in PNPC, and 49.86% in FOR. This indicates that Joint_CNN is more accurate in predicting clean modules, more effective in predicting defective instances on average, and better at reducing false negative instances across the target projects. The $Cohen^{\prime}s\ d$ results indicate a large effect size in F-measure and a moderate effect size in AUC, further confirming the superiority of Joint_CNN over Base_CNN. According to Table VIII, joint_CNN outperforms SVM, RF, DBN and AutoSpearman in terms of both F-measure and AUC. Specifically, Joint_CNN achieves average improvements of 42.86%, 30.63%, 61.75% and 2.13% in F-measure compared to SVM, RF, DBN and AutoSpearman, respectively. Similarly, in AUC, Joint_CNN demonstrates average improvements of 10.93%, 15%, 20.75% and 0.97%. The $Cohen^{\prime}s\ d$ results further validate the performance advantages of Joint_CNN. In terms of F-measure, the effect sizes are large, indicating significant improvements over SVM, RF and DBN. While the effect size for AutoSpearman is smaller, Joint_CNN still demonstrates overall better performance. For AUC, the effect sizes vary across different comparisons. Specifically, the effect size for SVM is medium, while the effect sizes for RF and DBN are large, indicating significant improvements. Although the effect size for AutoSpearman is small, Joint_CNN still demonstrates overall robustness.
TABLE VIII: The F-measure and AUC performance of different models on all datasets
F-measure AUC
Dataset SVM RF DBN AutoSpearman Joint_CNN SVM RF DBN AutoSpearman Joint_CNN
JDT 0.6239 0.6989 0.7162 0.7824 0.8544 0.7554 0.7419 0.7852 0.8598 0.8360
LC 0.3467 0.7177 0.3467 0.7871 0.8285 0.5846 0.8462 0.5846 0.5046 0.6862
ML 0.4919 0.5958 0.3219 0.8054 0.7123 0.6732 0.7522 0.4596 0.7339 0.6418
PDE 0.5565 0.6209 0.4874 0.7837 0.7111 0.6875 0.6601 0.6130 0.6957 0.7032
jm1 0.4476 0.5849 0.4476 0.7580 0.8621 0.5030 0.5762 0.5030 0.6269 0.6280
kc1 0.5695 0.5371 0.6918 0.5814 0.7303 0.6992 0.6141 0.7724 0.8111 0.8087
mc1 0.4754 0.5513 0.4742 0.5814 0.8621 0.8914 0.5577 0.8326 0.9373 0.6280
PC1 0.3870 0.7556 0.3939 0.8276 0.8621 0.6562 0.8557 0.6057 0.6815 0.6280
PC2 0.4736 0.4996 0.4213 0.9819 0.8621 0.4507 0.5000 0.3647 0.4194 0.6280
PC3 0.5580 0.5745 0.5224 0.8345 0.7957 0.7914 0.5772 0.7356 0.8336 0.8324
PC4 0.5700 0.6694 0.3535 0.8617 0.7472 0.7380 0.6154 0.5651 0.7814 0.7843
ant-1.7 0.7900 0.7121 0.5446 0.7760 0.7907 0.8409 0.7477 0.6591 0.8266 0.8022
camel-1.6 0.4920 0.5367 0.6314 0.6710 0.7337 0.6199 0.5621 0.6923 0.6356 0.6687
ivy-1.2 0.6244 0.7317 0.5000 0.8721 0.8347 0.6719 0.8125 0.7500 0.6250 0.8594
jedit-4.1 0.3914 0.3273 0.3431 0.7585 0.6946 0.4657 0.3743 0.4771 0.8698 0.8125
log4j-1.0 0.4286 0.3538 0.4286 0.6533 0.6520 0.6364 0.5909 0.6364 0.7750 0.8500
lucene-2.4 0.7153 0.3585 0.5983 0.7059 0.6947 0.7140 0.5000 0.6333 0.8250 0.7643
poi-3.0 0.6584 0.6786 0.5139 0.7819 0.6951 0.6463 0.6624 0.5392 0.8082 0.7414
synapse-1.2 0.5439 0.6067 0.3210 0.5247 0.6486 0.5437 0.6062 0.4125 0.6732 0.7516
xerces-1.3 0.6430 0.6865 0.4697 0.7611 0.8407 0.7747 0.6758 0.6465 0.7766 0.7875
Average 0.5394 0.5899 0.4764 0.7545 0.7706 0.6690 0.6453 0.6146 0.7350 0.7421
Improvement 42.86%↑ 30.63%↑ 61.75%↑ 2.13%↑ - 10.93%↑ 15.00%↑ 20.75%↑ 0.97%↑ -
$Cohen^{\prime}s\ d$ 2.3697 1.7196 2.9703 0.1712 - 0.7337 0.9543 1.1906 0.0657 -
We believe there are two reasons related to the improvement: (1) We explicitly integrate the prediction error $L_{pred}(f,D)$ into the loss function of the joint learning, enabling the predictor to capture discriminative information contained in the data during the training process, and subsequently improving the corresponding accuracy. (2) The $L_{int}$ can be regarded as a penalty term of $L$, which effectively encourages the framework to learn more generalizable representations of the data, leading to improved predictive accuracy for the data that is not contained in training dataset.
IV-C3 The results of RQ3
To demonstrate the global interpretability of our approach, we take the PC2 dataset as an example. As presented in Fig. 5, the decision logic can be abstracted as the path from the “DECISION_DENSITY” metric (denoted as the root) to the leaf node (denoted as 1 or 0). For example, one feasible decision path can be presented as DECISION_DENSITY -> MAINTENANCE_SEVERITY -> PERCENT_COMMENTS -> PERCENT_COMMENTS -> 0, which is similar to the form of decision tree’s interpretation. So, we can conclude our approach achieves the global interpretability.
We measure the computational cost by calculating the average training time. Specifically, we employ the Python $\mathrm{time}$ module to record the start and end times of the training process for both our approach and SP-Lime. To ensure statistical reliability, this procedure is repeated ten times for each method. The results show that the average computational costs for our approach and SP-Lime are 744.7 seconds and 2301.4 seconds, respectively. These findings indicate that our method achieves a significant reduction in computational cost compared to SP-Lime. The substantial difference can be attributed to the optimized design of our method, which abstracts the decision logic as the path of SDT from the root node to the leaf node. In contrast, SP-Lime requires extensive perturbations and evaluations across the feature space, leading to higher computational requirements.
Furthermore, we calculate the CD, FI and AI values for LC dataset to evaluate the reliability of our approach and SP-Lime. As presented in table IX, our method achieves a better performance in all three indicators, indicating our method is more reliable than SP-Lime. The reason for this result can be partially attributed to the random sampling involved in perturbation processes of SP-Lime [55].
TABLE IX: The CD, FI and AI values of Joint_SDT and SP-Lime
Method CD FI AI
Joint_SDT 1.00 0.8714 0.8429
SP-Lime 0.81 0.7933 0.8012
IV-C4 The results of RQ4
We conducted an ablation experiment on 5,589 instances in the PC2 dataset to analyze the impact of different features on model performance. The interpretations generated by our approach are presented in Fig. 5. In the interpretations of our approach, the metrics that have a greater impact on the prediction results are located closer to the root node.
Meanwhile, the impacts on the F-measure, AUC, and MCC after removing certain metrics are presented in Fig. 7. We can see that the accuracy of Joint_CNN went down significantly across MCC, F-measure and AUC three indicators after removing the “DECISION_DENSITY”. This indicates that this metric is critical for the accuracy of the predictor. This finding is consistent with the interpretation of our approach, as our approach placed it at the root node of SDT, indicating its decisive role in the model’s decision-making process. Thus, the significant decline in performance after removing this feature further corroborates its key role in the model. The importance of this feature is strongly evidenced both by the performance change and its position in the explanation.
To enhance the comprehensiveness of the analysis, the ablation experiment was extended to all datasets. The most important features interpreted by our approach for each dataset, as well as the ranking of performance degradation on the F-measure, AUC, and MCC after removing these features, are presented in Table X. The results demonstrate that the performance degradation is consistent with our explanatory results, further validating that our method effectively captures key features to improve prediction performance and provide accurate model interpretations.
TABLE X: Ranking of Feature Importance Based on SDT Interpretations and Ablation Results
Dataset Top Feature Identified by SDT F-measure Rank AUC Rank MCC Rank
JDT WCHU_numberOfMethodsInherited 4 9 5
LC WCHU_cbo 2 1 1
ML ck_oo_numberOfPublicMethods 5 9 5
PDE WCHU_fanOut 3 3 5
jm1 HALSTEAD_DIFFICULTY 7 7 1
kc1 v 1 2 2
mc1 LOC_TOTAL 1 1 1
PC1 branchCount 2 1 1
PC2 DECISION_DENSITY 1 4 1
PC3 DESIGN_DENSITY 4 3 3
PC4 LOC_TOTAL 2 2 1
ant-1.7 mfa 1 5 1
camel-1.6 lcom 4 3 4
ivy-1.2 cam 2 7 4
jedit-4.1 lcom3 6 10 4
log4j-1.0 max_cc 4 4 3
lucene-2.4 dam 1 1 1
poi-3.0 wmc 2 3 2
synapse-1.2 cbo 1 1 4
xerces-1.3 max_cc 2 8 2
(a) F-measure
(b) AUC
(c) MCC
Figure 7: Ablation study results across multiple indicators. The gray dashed line indicates the performance of DP-CNN before removing metrics.
V Threats to validity
While the proposed model demonstrates advancements in both prediction accuracy and interpretability, certain limitations may affect its generalizability and reliability. The details are as follows.
  • Choice of dataset: The experiment was conducted on the publicly available datasets. These datasets may not fully capture the diversity of software systems in real-world scenarios, potentially limiting the robustness of the model.
  • Hyperparameter setting: The model’s performance is related with the selection of hyperparameters for both the predictor and interpreter. Suboptimal tuning may lead to overfitting or underfitting, resulting in less accurate predictions or interpretations.
  • Application scenarios: In this paper, we focus on the within-project defect prediction that the training data and the test data are generated from the same project. However, the performance of our model in the context of the cross-project setting is not evaluated.
VI Conclusion and future works
In this paper, we explore the possibility of designing a defect predictor and its corresponding interpreter collaboratively. Unlike most existing approaches, which treat defect prediction and interpretation as separate tasks, we consider them strongly correlated. Based on this assumption, we introduce a framework aimed at improving the reliability of interpretation while simultaneously enhancing predictive accuracy by utilizing interpretation results. Based on the proposed framework, we conducted extensive empirical evaluations, which revealed the following key findings:
  • Compared to existing interpretation methods such as LIME and BreakDown, the proposed framework significantly improves the consistency and reliability of interpretations, providing robust explanations from a local perspective. This improvement is attributed to the incorporation of the Knowledge Distillation (KD) principle, which enables the accurate transfer of decision-making knowledge from the predictor to the interpreter.
  • The framework demonstrates superior performance on widely used datasets in terms of both F-measure and AUC, outperforming baseline models such as SVM, RF, DBN, and Base_CNN. By explicitly incorporating interpretation results into the loss function, the framework captures more discriminative information, leading to more robust predictive outcomes.
  • The proposed method offers global interpretability, providing developers with transparent global insights and decision-making rationales. This capability ensures that the framework not only achieves high accuracy but also delivers comprehensive explanatory support for real-world defect prediction tasks.
  • Ablation analysis further demonstrates that the framework effectively identifies key features that have the greatest impact on prediction outcomes, showcasing its ability to provide reliable explanations that align closely with the actual importance of these features.
Future research will investigate several possibilities to further improve the proposed model, including the following:
  • Expand the research by including a wider array of datasets that accurately reflect real-world software systems. Evaluate the robustness and generalizability of our model accordingly.
  • Incorporate automatic hyperparameter optimization methods into our model to reduce the likelihood of suboptimal tuning.
  • Assess the model’s performance in several defect prediction scenarios, such as the cross-project defect prediction.

### v2

@@
 Abstract
-Over the past fifty years, numerous software defect prediction (SDP) approaches have been proposed.
-However, the ability to explain why predictors make certain predictions remains limited.
-Explainable SDP has emerged as a promising solution by using explainable artificial intelligence (XAI) methods to clarify the decision-making processes of predictors.
-Despite this progress, there is still significant potential to enhance the reliability of existing approaches.
-To address this limitation, we treat defect prediction and the corresponding interpretation as two distinct but closely related tasks and propose a joint learning framework that allows for the simultaneous training of the predictor and its interpreter.
+Understanding why defect predictors classify software components as defective or clean is essential for software engineers that helps identify the root causes of defects and develop actionable bug-fixing plans.
+Existing solutions employ various explainable artificial intelligence (XAI) methods to clarify the decision-making processes of defect predictors.
+However, these post-hoc explanation techniques have two main limitations: 1.
+the interpretation results do not accurately reflect the model’s decision logic, and 2.
+they do not contribute to improving the performance of defect predictors.
+To address these limitations, we treat defect prediction and its corresponding interpretation as two distinct but closely related tasks, proposing a joint learning framework that trains the predictor and its interpreter simultaneously.
 The novelty of our approach lies in two main aspects: 1.
-We design feedback loops that convey the decision-making logic from the predictor to the interpreter.
-This ensures a high level of conciseness in decision logic and feature engineering for both the predictor and the interpreter, enabling the interpreter to achieve reliable local and global interpretability.
+we design a feedback loop that transfers decision logic from the predictor to the interpreter, ensuring a high degree of conciseness for both components.
 2.
-We incorporate the interpretation results as a penalty term in the loss function of the joint-learning framework.
-This not only improves the accuracy of the predictor but also imposes a stronger constraint on the reliability of the interpreter.
-We validated our proposed method against several existing explainable SDPs across multiple datasets.
-The results demonstrate its effectiveness in both interpretation and defect prediction.
-The source code for the proposed method is available at: https://github.com/BugPredictor/software-defect-prediction.git
+we incorporate interpretation results as a penalty term in the loss function of the joint learning framework.
+This not only enhances the accuracy of the predictor but also proposes a stronger constraint on the reliability of the interpreter.
+We validate our method against several existing explainable software defect predictors across multiple datasets.
+The results demonstrate its effectiveness.
+The source code of our method is available at: https://github.com/BugPredictor/software-defect-prediction.git
 Index Terms:
@@
 I Introduction
-For nearly half a century, software defect prediction (SDP) has been a significant research focus in the field of software engineering [1, 2].
-Its primary aim is to identify defective code early in the software development process, helping practitioners prioritize quality assurance tasks, especially when resources are limited or deadlines are tight [49].
-Inspired by the success of machine learning (ML) in various disciplines, such as computer vision and machine translation, researchers have explored applying different ML algorithms to improve the accuracy of SDP.
-However, despite numerous efforts, existing SDP models often fail to provide actionable guidance for practitioners, making practical application challenging [51].
-Recent studies have highlighted that the limited applicability of SDP in real-world scenarios is closely linked to the lack of interpretability [3, 4].
-Interpretability refers to the ability to provide comprehensible insights into the decision-making logic for humans.
-Without it, developers struggle to understand the logic behind SDP predictions, leading to distrust and reluctance to adopt these techniques [3, 5].
-Explaining prediction results can be particularly challenging for ML-based predictors, especially those using deep learning (DL).
-The complex structures and large number of parameters in such models obscure straightforward interpretation.
-To address this issue, researchers have investigated the feasibility of explainable SDP, incorporating explainable artificial intelligence (XAI) into SDP.
-Efforts in explainable SDP are generally categorized into two groups [6]: global and local interpretations.
+Software defect prediction (SDP) has attracted significant attention in software engineering for over five decades [1, 2].
+The primary goal of this research is to identify defective code early in the software development process, helping practitioners prioritize quality assurance tasks, especially when resources are limited or deadlines are tight [3].
+In recent years, inspired by the success of machine learning (ML) and deep learning (DL) in various disciplines, researchers have introduced different ML/DL algorithms into SDP.
+Despite significant improvements in the performance of SDP models, many practitioners remain hesitant to integrate SDP techniques into practical applications [4].
+Such a situation is primarily due to the lack of interpretability of these models.
+In practical scenarios, practitioners often ask: What is the reason for predicting a code fragment as buggy?
+How do different metrics influence the prediction?
+These questions underscore an urgent requirement for explainability in SDP [5, 6, 7].
+To address the interpretation issue, researchers have proposed various explainable SDP models that integrate different explainable artificial intelligence (XAI) techniques with SDP.
+Existing explainable SDP models are generally categorized into two groups [8]: global and local approaches.
 Global approaches focus on explaining the overall behavior of a defect predictor, offering insights into how the model makes decisions across all instances [9].
-Local approaches provide clear evidence for why a given instance is predicted as buggy or clean.
-The prediction results provide actionable guidance for practitioners to address key questions, such as why a specific prediction was made and how to mitigate associated risks [57].
-As a result, practitioners often prefer to use local interpretation techniques in practice [3].
-Nevertheless, local techniques often encounter challenges with reliability under different settings [52].
+In contrast, local approaches provide clear evidence for why a given instance is predicted as buggy or clean.
+Compared to global approaches, local approaches focus on identifying the root cause of specific defects and provide actionable guidance for practitioners to address key questions, such as why a specific prediction was made and how to mitigate associated risks [10].
+As a result, local interpretation techniques are generally considered more favorable in practice [5, 11].
+However, the task properties of interpretation and defect prediction are quite different, making explainable SDP a challenging problem.
+On one hand, SDP relies on high-complexity ML/DL models to capture intricate relationships among software metrics and defect patterns.
+These models are often considered as black boxes, with decision-making processes that are difficult to interpret.
+On the other hand, XAI methods favor simpler models that provide a clear, human-understandable logical reasoning path.
+Bridging these two competing objectives—complexity and interpretability—remains a significant challenge.
+Most existing solutions adopt a post-hoc explanation strategy, in which an SDP model is first trained and subsequently interpreted by XAI techniques [12, 13].
+Although this strategy offers a degree of interpretability, it presents two main challenges: 1.
+the interpretations often fail to reliably reflect the model’s actual decision logic, and 2.
+they do not contribute to improving the predictive performance of defect models.
 [The interpretation for the SMOTE-based predictor] Refer to caption [The interpretation for the RUS-based predictor] Refer to caption
@@
 (b) (b) Two interpretations produced by LIME
-Figure 1: Interpretations of LIME for a code fragment.
-To illustrate the challenges associated with local approaches, we take the interpretation for ”AbstractCommand.java” from the Apache ActiveMQ project as an example.
-The explanations generated by LIME with different samplers (i.e., SMOTE and RUS) are shown in Fig.
-1 LABEL:sub@fig_1:LIME_RUS.
-The explanations generated by applying LIME twice to the same file are shown in Fig.
+Figure 1: An example of unreliable interpretations.
+To better illustrate the challenges associated with local approaches, we examine the interpretation of ”AbstractCommand.java” from the Apache ActiveMQ project as an example.
+The explanations generated by LIME using different data sampling techniques (i.e., SMOTE and RUS) are shown in Fig.
+1 LABEL:sub@fig_1:LIME_RUS, and the explanations generated by executing LIME twice are shown in Fig.
 1 LABEL:sub@fig1:LIME_two.
-These figures present the rankings of features that contribute to the predictions.
-Features displayed in red indicate their contribution to predicting buggy code, while those displayed in green indicate their contribution to the prediction of clean code.
-In Fig.
-1 LABEL:sub@fig_1:LIME_RUS, we can see that although the SDP model with different samplers predicts the file as buggy, their explanations vary significantly.
-Among the ten features selected by LIME for the SMOTE-based predictor, only two are also selected for the RUS-based predictor.
+The features highlighted in red contribute to the prediction of buggy code, while those highlighted in green contribute to the prediction of clean code.
+From the Fig.
+1 LABEL:sub@fig_1:LIME_RUS, we observe that the explanations vary significantly.
+Among the ten features selected by SMOTE-based LIME, only two are also selected by RUS-based LIME.
 Fig.
-1 LABEL:sub@fig1:LIME_two further demonstrates that the two explanations generated by running the same predictor twice show considerable variation.
-For instance, in the first explanation, seven features contribute to the prediction of buggy code, whereas in the second explanation, the number increases to eight.
-To address this issue, we consider defect prediction and the corresponding interpretation as two distinct but strongly related tasks and propose a joint-learning framework in this paper.
+1 LABEL:sub@fig1:LIME_two further illustrates the issue of poor reliability.
+In the first round of explanation, seven features are identified as correlated with buggy code, whereas in the second round, this number increases to eight.
+To address this issue, we treat defect prediction and the corresponding interpretation as two distinct yet closely related tasks, proposing a joint learning framework in this paper.
 Specifically, we design multiple feedback loops between the interpreter and predictor to simultaneously achieve reliable interpretation and high prediction accuracy.
-The specific contributions are as follows:
-  • To improve the reliability of interpretation, we revised the loss function of the knowledge distillation (KD) framework and introduced output fidelity loss and feature fidelity loss.
-This strategy constructs a knowledge transfer loop between the interpreter and predictor, ensuring a high level of conciseness in decision-making logic and feature engineering power.
-  • To improve the accuracy of the predictor, we incorporate the interpretation results as a penalty term within the loss function of the joint-learning framework.
-This enables the predictor to capture more discriminative features guided by the interpretation results while imposing stronger restrictions on the fidelity of the interpreter to the predictor.
-  • Extensive experiments are conducted on 20 datasets with an in-depth analysis.
-The results indicate that the joint-learning framework improves both the reliability of interpretation and the accuracy of the predictor.
-The paper is organized as follows: Section II provides an overview of the existing research on SDP and interpretation methods used in SDP.
-In Section III, we describe our proposed method in detail, while in Section IV, we provide the details of the experiment.
-In Section V,we discuss the threats to validity.
-Finally, in Section VI, we present our conclusion and future works.
+The main contributions are as follows:
+  • To enhance the reliability of interpretation, we introduce two novel loss functions: output fidelity loss and feature fidelity loss.
+This strategy establishes a knowledge transfer loop between the interpreter and predictor, ensuring both a high level of conciseness in decision-making logic and feature engineering power between these two components.
+  • To improve the accuracy of the predictor, we incorporate the interpretation results as a penalty term in the loss function of the joint learning framework.
+This strategy enables the predictor to capture more discriminative features guided by the interpretation results while simultaneously imposing stronger constraints on the fidelity of the interpreter to the predictor.
+  • Extensive experiments are conducted on twenty datasets with an in-depth analysis.
+The results demonstrate that the joint learning framework improves both the reliability of interpretation and the accuracy of the predictor.
+The remainder of paper is organized as follows: Section II provides an overview of the existing research on SDP and interpretation methods used in SDP.
+Section III describes our proposed method in detail, while Section IV presents the experimental setup and results.
+Section V discusses the threats to validity.
+Finally, Section VI presents our conclusions and outlines directions for future work.
 II Related Work
-This section provides an overview of SDP and explores existing techniques for explainable SDP.
 II-A Software defect prediction
-The primary goal of SDP is to classify code fragments (e.g., method or class) as either clean or buggy.
-Constructing a typical SDP model involves three steps.
-The first step is to transform software metrics, such as cohesion, coupling, complexity, encapsulation, inheritance, and size, into corresponding numerical representations [10].
+Constructing an SDP model involves three steps.
+The first step is to transform software metrics, such as cohesion, coupling, complexity, into numerical representations [14].
 The second step labels code fragments as buggy or clean based on post-release defects.
-These defects can be collected from a Bug Tracking System (BTS) by connecting bug reports to their bug-fixing activities.
-Code fragments associated with these bug-fixing activities are regarded as buggy, while all others are labeled as clean.
-Finally, the labeled code fragments serve as training data for an ML-based classifier, which is then used to predict whether code fragments in a later version are buggy or clean.
-Numerous ML algorithms, including Fuzzy self-organizing maps, K-means, SVM, random forest, logistic regression, etc., have been extensively employed in previous research [15, 16, 17, 18, 19, 20, 21].
-However, metrics used in SDP consistently exhibit correlations and redundancies [11], compromising the effectiveness of traditional ML-based predictors [12].
-To address this issue, various DL models have been investigated to improve SDP performance.
-The initial DL-based defect predictor, Deeper, introduced by Yang et al.
-[22], employed a deep belief network to identify discriminative information in the input data.
-Experimental results showed that Deeper could identify 32.22% more defects than many traditional ML-based approaches.
-Inspired by Yang’s work, many researchers have proposed different DL-based solutions.
-Specifically, Hoang et al.
-proposed a convolutional neural network (CNN)-based defect prediction model [23], which experimental results indicated outperformed Deeper.
-Qiao et al.
-proposed another DL-based model called DPNN to predict the number of defects [24].
-Experimental results showed a 14% reduction in mean square error and an 8% increase in the squared correlation coefficient.
-Li et al.
-proposed another CNN-based defect predictor, DP-CNN [25], which achieved better performance than other methods.
-As demonstrated, DL has significantly improved the performance of SDP.
-To ensure the generalization, we employ the DP-CNN proposed in [25] as the predictor and evaluate the difference in accuracy before and after integrating it into the joint-learning framework.
-II-B Interpretation of defect prediction
-In many practical application scenarios, practitioners often ask: What is the underlying decision logic of the predictor?
-What is the reason for predicting a code fragment as buggy?
-How do different metrics influence the prediction?
-These questions highlight an urgent requirement for explainable defect prediction.
-However, highly accurate SDP models (e.g., Random Forest, DL) are often black-box in nature, making them difficult to interpret.
-Efforts to improve the interpretability of predictions can be broadly divided into two main categories [6]: global and local approaches.
-The global approaches offer a high-level view of how inputs influence the model’s predictions and reveal patterns that are consistent across the model’s entire operation, rather than focusing on individual predictions.
-Many ML algorithms (e.g., decision trees, logistic regression techniques) and statistical methods (e.g., ANOVA, variable importance) fall into this category[26, 27].
-These methods share two common aspects: 1.
-Simulatability: Every step of the decision-making process is comprehensible.
-For example, decision trees are simulatable since we can observe the state of every node as well as the feature thresholds used to separate samples.
-2.
-Transparency: The decision-making process can be broken down into interpretable steps.
-For instance, the interpretability of a decision tree is defined by the computational path from the root node to the leaf nodes.
-However, global approaches cannot provide detailed interpretations for each code fragment, and their relatively simple structure often fails to guarantee optimal predictive accuracy.
-Consequently, many practical applications have adopted local methods.
-In contrast to global approaches, local methods such as LIME [8] and BreakDown [28], usually interpret the prediction of each code fragment as a list of ordered features.
-Feedback from practitioners indicates that the interpretations produced by local methods can offer more actionable guidance [52].
-However, recent studies have reported that local techniques lack reliability under various conditions [30, 31, 52].
-Firstly, the interpretations of the same code can be inconsistent when different data sampling techniques are used.
-In SDP, sampling is frequently used due to the scarcity of buggy codes.
-It is an effective way to mitigate the effects of imbalanced data on the ML algorithms [53].
-In [52], Jiho Shin et al., investigated the consistency of local methods across five commonly used data sampling techniques.
-Their analysis revealed that both the features and the ranking of features in the interpretation vary significantly when different data sampling techniques are employed.
-Secondly, the interpretations of a code fragment do not exhibit consistency when multiple interpretations are executed.
-The underlying intuition of local methods regards the interpreter $g$ and the predictor $f$ as two independent ML algorithms and leverages $g$ to approximate the behavior of $f$ within the local area of an instance $x$, denoted as $\pi_{x}$.
-In general, local methods cannot guarantee that the $\pi_{x}$ required for multiple interpretations of the same instance are the same.
-So different $\pi_{x}$ can lead to distinct interpretations for the same code fragment [31].
-Finally, the predictions produced by the interpreter and the predictor for the same code fragment are not consistent.
-To reduce the difficulty of interpretation, many local methods define the interpreter as a simple-structure ML model.
-Taking the most widely used local interpretation technique LIME as an example, it sets a linear regression model as the interpreter, and the interpretation of SDP is oversimplified as a linear transformation between input metrics and prediction outcomes.
-Recently, [31] emphasized that attempting to comprehend an intricate model by employing a simple model might be overly optimistic.
-The inconsistency between the predictions generated by the interpreter and the predictor indicates that the interpretation fails to capture the underlying decision-making logic of the predictor.
+Finally, leverage the labeled code fragments to train an ML/DL-based classifier for identifying future defect-prone code.
+Over the past fifty years, numerous ML algorithms, including fuzzy self-organizing maps, K-means, SVM, Random Forest, and logistic regression, have been extensively employed in SDP research [15, 16, 17, 18, 19, 20, 21].
+However, software metrics often exhibit correlations and redundancies [22], and traditional ML algorithms struggle to capture the complex dependencies among them [23].
+To address this issue, various DL models have been introduced in SDP.
+One of the earliest explorations, Deeper [24], employed a deep belief network to identify discriminative information in input data.
+Experimental results showed that Deeper identified 32.22% more defects than many traditional ML-based approaches.
+Inspired by the Deeper, researchers have introduced a sophisticated DL algorithms into SDP to enhance performance and broaden application scenarios.
+For example, [25, 26] leaverage convolutional neural networks (CNNs) to build defect prediction models, achieving better performance than other DL-based approaches.
+Additionally, Qiao et al.
+employed a DL model to predict the number of defects [27].
+Experimental results showed a 14% reduction in mean squared error and an 8% increase in squared correlation coefficient.
+To ensure generalizability, we employ DP-CNN, proposed in [26], as the predictor and evaluate the accuracy differences before and after integrating it into the joint learning framework.
+II-B Explainable defect prediction
+Efforts in explainable SDP can be broadly divided into two main categories [8]: global and local approaches.
+The global approaches provide a high-level view of how inputs features influence the model’s predictions.
+Many ML algorithms (e.g., decision trees, logistic regression) and statistical methods (e.g., ANOVA, variable importance) fall into this category[28, 29].
+However, global approaches cannot offer detailed interpretations for individual code fragment, and their simple structure often fails to ensure optimal predictive accuracy.
+As a result, many practical applications have adopted local methods instead.
+Local interpretation methods such as LIME [12] and BreakDown [30], primarily focus on interpreting specific code fragments.
+However, recent studies have reported that local techniques often lack reliability under various conditions [31, 32, 11].
+Firstly, interpretations can be inconsistent when different data sampling techniques are used or when the same interpreter is executed multiple times.
+Local interpretation methods typically employ an interpreter to approximate the predictor’s behavior within a local area of the given instance.
+The local area is generated by the sampling algorithm.
+However, the randomness of the sampling process makes it difficult to maintain consistency across local areas—regardless of whether the same or different sampling methods are used [11].
+As a result, the same code fragment may yield distinct interpretations[32].
+Secondly, interpretations tend to be oversimplified.
+Local interpretation methods usually employ ML algorithms with simple structure as interpreters.
+For example, LIME utilizes a linear regression model as the interpreter.
+The interpretation of SDP is oversimplified as a linear transformation between input metrics and prediction outcomes.
+Recent study [32] reported that attempting to comprehend an intricate model by employing a simple model might be overly optimistic.
+Such interpretations often fail to capture the underlying decision-making logic of the predictor.
 II-C Knowledge distillation
-Knowledge distillation (KD) is a typical model compression technique that transfers knowledge from a teacher model (e.g., DL models) to a student model (e.g., a shallow neural network) [32, 33].
-KD is regarded as a promising framework for improving the reliability of interpretation due to the faithful knowledge transfer channel between the student and teacher models.
-In the health care domain, [34] leveraged distilled knowledge to discover disease patterns.
-Similarly, in the business domain, [35] distilled dynamical pricing knowledge from a complex black-box DL model, optimizing revenue while maintaining interpretability.
-In this paper, we designate the predictor as the teacher model and the interpreter as the student model, and transfer the decision-making knowledge of predictor to interpreter based on the KD principle.
-The main difference between our approach and existing methods lies in the revised loss functions of KD framework, which enable us to train the interpreter and predictor in a collaborative manner.
+Knowledge distillation (KD) is a model compression technique that transfers knowledge from a teacher model (e.g., DL models) to a student model (e.g., a shallow neural network) [33, 34].
+It is considered a viable solution for enhancing the reliability of interpretation due to the knowledge transfer mechanism between the teacher and student models.
+For instance, [35] utilized distilled knowledge to identify disease patterns, while [36] distilled dynamic pricing knowledge from a complex black-box DL model.
+In this paper, we treat the predictor and the interpreter as the teacher and student models, respectively, and design knowledge transfer channels based on the KD principle.
+The key distinction of our approach lies in the collaborative training mechanism: unlike existing KD methods, which keep the parameters of the teacher model fixed, our method trains both the interpreter and the predictor in a collaborative manner.
 III Proposed Method
-Fig. 2 provides an overview of our approach, which consists of four modules: metric selection, predictor and interpreter design, joint learning framework, and local interpretation.
+Fig. 2 provides an overview of our proposed approach, which consists of four modules: metric selection, predictor and interpreter design, joint learning framework, and local interpretation.
 Figure 2: The framework of our approach.
 III-A Metric selection
-We carefully selected 16 metrics for our method after a thorough investigation of existing literature and metric suites, including the works of [10, 36, 37].
-The details of metrics are presented in Table I.
-The selection of these metrics was guided by two primary criterias, ensuring their relevance and effectiveness in the context of defect prediction: (1) Expressiveness: The metrics needed to be sufficiently expressive, enabling them to capture subtle variations in code quality, structure, and complexity between similar pieces of source code.
-(2) Proven Effectiveness: Each selected metric has been validated in prior studies [17, 22, 38, 39], demonstrating its utility in software defect prediction tasks.
-By adhering to these criterias, we constructed a dataset that covered a broad spectrum of code attributes, ranging from complexity measures to process-related characteristics.
+We selected sixteen metrics for our method based on a review of existing literature [14, 37, 38].
+The details of these metrics are presented in Table I.
+The selection was guided by two criterias: (1) expressiveness: The metrics needed to be sufficiently expressive to capture subtle variations in code quality, structure, and complexity.
+(2) effectiveness: Each selected metric has been demonstrated its utility in SDP tasks [17, 24, 39, 40].
+According to these criteria, we constructed a dataset that encompasses a broad spectrum of code attributes, ranging from complexity measures to process-related characteristics.
 TABLE I: Metrics suite
@@
 LOC Lines of Code
-CC McCabe’s cyclomatic complexity
+$C$ McCabe’s cyclomatic complexity
 III-B Defect prediction and interpretation model
@@
 III-B1 Defect predictor
-To ensure the generality of our method, we chose a DL-based model, DP-CNN [25] as the predictor.
-It consists of three main components: the convolution layer, max pooling layer, and classification layer.
-To further improve accuracy, we added an attention layer to the original model.
-The overall architecture of DP-CNN is presented in Fig.
+In this paper, we chose a DL-based model, DP-CNN [26] as the predictor.
+It consists of three main components: a convolution layer, a max pooling layer, and a classification layer.
+To further improve predictive accuracy, we extend the original model by integrating an attention mechanism.
+The overall architecture of DP-CNN is illustrated in Fig.
 3.
 Figure 3: Model Architecture of DP-CNN.
-The working procedure of DP-CNN can be summarized as follows: First, the input defect metrics are fed into the convolution layer to extract correlations between features.
-Then feature maps are then passed through the max pooling layer, which reduces the dimensionality while retaining the most significant features.
-To further enhance the model’s ability to focus on critical patterns, the attention layer dynamically assigns weights to the feature maps.
-Finally, the processed features are passed through a fully connected neural network and a softmax classifier to generate the prediction result.
-The correlations between metrics are crucial for defect prediction.
-Two characteristics of the convolution layer, sparse connectivity and shared weights, make it particularly suitable for this task [25].
-Sparse connectivity enables CNNs to have a wider receptive field, capturing non-linear correlations between metrics.
-Shared weights ensure that each CNN kernel has the same parameters, allowing predictors to capture these correlations regardless of their position in the input metrics.
-The max pooling layer retains the most significant features by selecting the maximum local element in the convolutional output, enhancing the model’s robustness to changes in the position of metrics in the input.
-The attention layer dynamically assigns weights to the inputs, providing a flexible way to capture correlations between feature maps.
-The fully connected neural network is responsible for converting the feature maps generated by previous layers into a logit.
-The softmax block is positioned at the end of the entire neural network, transforming the logits into the probabilities.
+The DP-CNN model predicts defect-proneness through the following process: Selected metrics are first fed into the convolutional layer to capture correlations among features.
+The output is then passed through the max pooling layer to highlight key features while reducing redundancy.
+Next, the attention layer focus on critical defect patterns by dynamically assigning weights.
+Finally, a fully connected network followed by softmax classifier generates the prediction.
+We choose the CNN-based model as the predictor for three key reasons: (1) Sparse connectivity enables a wider receptive field to capture non-linear correlations.
+(2) Weight sharing ensures positional invariance in feature extraction [26].
+(3) Max pooling enhances robustness by preserving significant features, while the attention layer further refines feature importance.
 III-B2 Interpretation model
@@
 The SDT is a variant of the fuzzy decision tree, sharing a similar structure with the ordinary decision tree but differing in the definition of nodes.
-In SDT, nodes are defined as neurons with learnable weight $W$ and bias $b$.
-Weights reflect the importance of each feature in the decision-making process at the node.
+In SDT, nodes are defined as neurons with learnable weights $W$ and biases $b$.
 The output of each node is defined as $P_{i}(x_{i})=\sigma(x_{i}W_{i}+b_{i})$, which determines the probability of transitioning to the right subtree.
 The $\sigma(\cdot)$ denotes the activation function and $x_{i}$ is the input of the $i$-th node.
-The output at the $l$-th leaf is defined as:
+The output at the $l$-th leaf node is defined as:
 $Q_{k}^{l}=\frac{{\rm exp}(\phi_{k}^{l})}{\sum_{k}{\rm exp}(\phi_{k}^{l})}$ (1)
 where $Q_{k}^{l}$ is the probability at the $l$-th leaf of type $k$ defect, and $\phi_{k}^{l}$ is the learned feature at that leaf.
-The reasons we use SDT as an interpreter are as follows: Owing to its neural-like node architecture, SDT exhibits comparable feature engineering capabilities to CNN, enabling it to accurately retain the decision-making knowledge of CNN.
-Furthermore, in comparison to CNN, the complexity of SDT has been reduced, rendering the decision-making process more comprehensible.
+The reasons we use SDT as an interpreter are as follows: Owing to its neural-like node structure, SDT exhibits comparable feature engineering capabilities to CNN, enabling it to simulate the decision-making knowledge of CNN.
+Furthermore, the complexity of SDT has been reduced, making the decision-making process more comprehensible.
 III-C Joint learning framework
 The architecture of the joint learning framework is presented in Fig.
-4, each feedback loop corresponds to a specific loss function.
-Training stops when the model reaches the maximum number of iterations or meets the early stopping condition.
-Given training data set $D=\{\left(x_{i},y_{i}\right)_{i=1}^{N}\}$, $x_{i}\in R^{n}$, $y_{i}\in\{0,1\}$, the predictor $f$ and interpreter $g$, the loss function of joint learning framework is formalized as follows:
+4, each feedback loop corresponds to a loss function.
+Given training data set $D=\{\left(x_{i},y_{i}\right)_{i=1}^{N}\}$, $x_{i}\in R^{n}$, $y_{i}\in\{0,1\}$, the predictor $f$ and interpreter $g$, the loss function for joint learning framework is defined as follows:
 Figure 4: The architecture of joint learning framework.
 $L=\alpha L_{pred}(f,D)+\beta L_{pred}(g,D)+L_{int}(f,g,D)$ (2)
-where $\alpha$ and $\beta$ are hyperparameters.
-The original loss function of KD consists of two essential parts: the distillation loss $L_{soft}$ (which is equivalent to $L_{int}(f,g,D)$ in Eq.
+where the hyperparameters $\alpha$ and $\beta$ denote the relative importances of the subloss to the the overall loss.
+The original loss function of KD consists of two parts: the distillation loss $L_{soft}$ (which is equivalent to $L_{int}(f,g,D)$ in Eq.
 (2)) and the student loss $L_{hard}$ (which is equivalent to $L_{pred}(g,D)$ in Eq.
-(2))).
+(2)).
 $L_{soft}$ quantifies the difference between the teacher model and the student model, which serves as a feedback loop for transferring knowledge from the teacher model to the student model.
-Minimizing $L_{soft}$ will enforce the student model to adopt similar decision logic to the teacher model.
 $L_{hard}$ evaluates the difference between the student model’s predictions and the ground truth.
-Minimizing $L_{hard}$ ensures that the student model retains the appropriate decision logic of the teacher model.
-In this paper, we introduce two key modifications to the existing loss function: (1) added the output fidelity loss $L_{pred}(f,D)$, and (2) revised the $L_{soft}$.
+In this paper, we introduce two key modifications to the original KD loss function: (1) added the output fidelity loss $L_{pred}(f,D)$, and (2) revised the $L_{soft}$.
 $L_{pred}(f,D)$ shares a similar definition with $L_{pred}(g,D)$ indicating the consistency of the predictor’s output to the ground truth.
@@
 $L_{pred}(g,D)=\sum_{i=1}^{|D|}\left(g(x_{i})-y_{i}\right)^{2}$ (4)
-We introduce output fidelity loss guided by two critical considerations: (1) Optimizing the predictor and interpreter simultaneously.
-Unlike the original KD framework, which keeps the parameters of teacher model (predictor) fixed during the distillation process, minimizing the output fidelity loss will enforce the predictor to capture the discriminative features of the input metrics.
-(2) Improving the fidelity of the interpreter.
+Introducing output fidelity loss here can enhance both prediction accuracy and interpreter fidelity simultaneously.
 Within the KD framework, predictor and interpreter are mutually dependent.
-Improving the accuracy of the predictor is equivalent to indirectly enforcing the interpreter to provide a more accurate explanation of the predictor’s decision logic.
-Therefore, introducing the output fidelity loss can boost the performance of the predictor and provide a more robust and insightful knowledge based on the KD principle.
-Moreover, we revised the loss function $L_{soft}$ (denoted as $L_{int}(f,g,D)$ in Eq.
-(2)).
-According to Eq.
-(5), the $L_{int}(f,g,D)$ involves two parts, $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$.
+Minimizing the output fidelity loss will encourage the predictor to achieve heigher accuracy.
+Moreover, improving the accuracy of the predictor is equivalent to indirectly enforcing the interpreter to provide a more accurate explanation of the predictor’s decision logic.
+Therefore, incorporating the output fidelity loss can boost the performance of the predictor but also yield a more robust and insightful knowledge based on the KD principle.
+The revised the loss function $L_{soft}$ (denoted as $L_{int}(f,g,D)$ in Eq.
+(2)) is presented in Eq.
+(5).
+It involves two parts, $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$.
 $L_{of}(f,g,D)$ is equivalent to the distillation loss in the original KD framework.
 The feature fidelity loss $L_{fef}(f,g,D)$ evaluates the similarity between the feature maps generated by predictor and interpreter.
-Minimizing $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$ can enforce the interpreter to adopt similar decision logic and feature engineering power of predictor.
+Minimizing $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$ can enforce the interpreter to achieve similar feature engineering power of predictor.
 $L_{int}(f,g,D)=\lambda L_{of}(f,g,D)+\gamma L_{fef}(f,g,D)$ (5)
-where $\lambda$ and $\gamma$ are hyperparameters. $L_{of}(f,g,D)$ shares a similar definition with $L_{soft}(f,g,D)$, which ensures the output of $g$ should be similar to that of $f$.
+where $\lambda$ and $\gamma$ are hyperparameters. $L_{of}(f,g,D)$ shares a similar definition with $L_{soft}(f,g,D)$.
 $L_{of}(f,g,T)=-\sum_{i=1}^{|D|}q_{i}^{T}\log(p_{i}^{T})$ (6)
@@
 where function $d(\cdot)$ adjusts $z_{i}$ to the same dimension of $v_{i}$.
-To further clarify the proposed joint learning framework, we take the PC2 dataset as an example.
-The data in PC2 involves software metrics and labels.
-Software metrics serve as the input for both the predictor $f$ and interpreter $g$.
-Labels indicate whether a code fragment is buggy or clean.
-The predictor and interpreter were optimized jointly using the loss function defined in Eq.
-(2).
-Through the joint training process, the predictor was continuously refined to improve its classification accuracy.
-This process enables the predictor to effectively distinguish between buggy and clean code fragments.
-Simultaneously, the interpreter was trained to generate a decision tree that aligns with the predictor’s decision-making logic.
-Through this process, the predictor was also influenced by the interpreter, as minimizing $L_{int}(f,g,D)$ required the predictor to refine its feature extraction and decision logic to better align with the explanations provided by the interpreter.
-In Fig.
-5, the decision tree generated by the interpreter reflects the optimized decision logic and highlights key decision paths for defect prediction.
-Figure 5: The interpretations of the proposed method on PC2 dataset.
+To further demonstrate the proposed joint learning framework, we use the PC2 dataset as an example.
+Each data point in PC2 consists of twenty software metrics and one label.
+According to Eqs.
+(3, 4), DP-CNN and SDT iteratively optimize their parameters to minimize discrepancies between predictions and actual labels.
+Crucially, improving prediction accuracy of DP-CNN inherently enforces the SDT to provide a more precise explanation of the predictor’s decision-making process, which indirectly improve interpreter’s reliability.
+Meanwhile, according to Eq.
+(6), the features generated by DP-CNN and SDT should be similar, ensuring that both models have comparable decision logic, thereby enhancing the interpretability of SDT.
+Additionally, Eq.
+(7) ensures consistency between SDT’s selected features and the input metrics, which strengthens the reliability of the interpreter.
+As presented in Fig.
+5, the decision logic of DP-CNN is interpretered as a set of “IF-THEN” paths in the SDT, which is similar to the decision tree doess [41].
+Figure 5: The interpretations of the proposed method on PC1 dataset.
 III-D Local interpretation
-It is obvious that SDT obtains the global interpretability as decision tree does [41], and interpretation for a dataset can be abstracted as a set of “IF-THEN” rules.
-In this section, we focus on the interpretability for an instance.
-Given a metric set $A=\{a_{1},a_{2},...a_{d}\}$, the sensitivity of SDT for an instance $x$ on the $a_{i}$ is defined as follows.
+In this section, we focus on the interpretability for an instance. Given a metric set $A=\{a_{1},a_{2},...a_{d}\}$, the sensitivity of SDT for an instance $x$ on the $a_{i}$ is defined as follows.
 $S(a_{i})=\frac{\Delta g_{i}(x)}{\Delta a_{i}}$ (8)
@@
 $\bar{a}_{i}=\frac{1}{N}\sum_{j}^{N}a_{j}^{i}$ (10)
-where $N$ is the number of training data, and $a_{j}^{i}$ is the $i$-the metric of $x_{j}$.
+where $N$ is the number of training data, and $a_{j}^{i}$ is the $i$-th metric of $x_{j}$.
 According to Eq.
-(8), we interpret the correlation between the instance and predictive result as the sensitivity of feature perturbation.
-Given an instance $x_{i}$, if a metric has a greater sensitivity, that means such a metric has a stronger influence on the prediction result.
-Since the $\Delta a_{i}$ is a constant and the parameters of interpreter are fixed after training, the sensitivity $S$ can provide a stable interpretation for any instance.
-The algorithm for local interpretation is presented below:
+(8), we interpret the correlation between the instance and predictive result as the sensitivity to feature perturbation.
+Given an instance $x_{i}$, if a metric has a greater sensitivity, this means the metric has a stronger influence on the prediction result.
+Since the $\Delta a_{i}$ is a constant and the parameters of interpreter are fixed after training, the sensitivity $S$ can provide a stable and consistent interpretation for any instance.
+The algorithm for performing local interpretation is presented below:
 Algorithm 1 Local Interpretation
 0: interpreter $g$, instance $x$, metric set $A=\{a_{1},a_{2},\dots,a_{d}\}$, and hyperparameters: $\lambda,\gamma$.
-0:  Sensitivity index $SI$
+0: Sensitivity index SISI
 1:for $a_{i}\in A$ do
-2:   Calculate $\text{Calculate }\bar{a}_{i}$ Eq. (10)
-3:   Calculate $\text{Calculate }\Delta a_{i}$ Eq. (9)
-4:   Calculate $\text{Calculate }\Delta g_{i}(x)$
+2:  $\text{Calculate }\bar{a}_{i}$ Eq. (10)
+3:  $\text{Calculate }\Delta a_{i}$ Eq. (9)
+4:  $\text{Calculate }\Delta g_{i}(x)$
 5:  $S=\cup\ S(a_{i})$ Eq. (8)
 6:end for
-7:  $SI\leftarrow$ Sorts $S$ in descending order
-8:  return $SI$
-The loop structure (lines 2 to 5 in the algorithm) calculates the sensitivity of each metric and generates the sensitivity set $S$.
+[…]
+8:return SISI
+The loop structure (lines 2 to 5 in the algorithm) computes the sensitivity of each metric and generates the sensitivity set $S$.
 Line 7 rearranges the sensitivity set $S$ in descending order and puts the index of $S$ to $SI$.
 IV Experiment
-To objectively evaluate the performance of our method, we conducted an experiment to quantify the performance of the proposed method and the baselines on 20 datasets, aiming to answer the following research questions:
-RQ1: What reliability-related improvements does the proposed method make in terms of consistency and faithfulness?
-As mentioned in Section II, many local methods often yield inconsistent interpretations by adopting different data sampling techniques as well as interpreting the same code fragment multiple times.
-These two factors greatly affect the reliability of explainable SDP.
+To objectively evaluate the performance of our method, we conducted an experiment to quantify the performance of the proposed method and the baselines on twenty datasets, aiming to answer the following research questions:
+RQ1: What improvements does the proposed method offer for the consistency of interpretations?
+As mentioned in Section II, many local methods tend to yield inconsistent interpretations, which greatly affects the reliability of explainable SDP.
 To evaluate the consistency of interpretation, we adopt the Coincidence Degree (CD) in RQ1 [42].
@@
 A greater $CD$ indicates greater consistency.
-Furthermore, as discussed in Section II, another issue related to the reliability of interpretation is that predictor and interpreter may generate different predictive results for the same instance.
+Furthermore, another issue related to the consistency of interpretation is that the predictor and interpreter may generate different predictive results for the same instance.
 To evaluate the consistency between predictor and interpreter, we introduce Fidelity of Interpretation (FI) [42, 43, 44] here.
@@
 $\mathbb{I}(\cdot)$ is an indicator function, which generates 1 when $f(x_{i})=g(x_{i})$, 0 otherwise.
-In addition to the FI, we argue that the prediction produced by an interpreter should also align with the ground truth.
-This alignment indicates how much correct decision knowledge is obtained from the predictive model.
-In this paper, we use the Accuracy of Interpreter (AI) as another indicator to measure this alignment.
+In addition to FI, we also argue that the prediction produced by an interpreter should also align with the ground truth.
+The alignment indicates how much correct decision knowledge is captured from the predictor.
+In this paper, we use the Accuracy of Interpreter (AI) to measure this alignment.
 Given the dataset $D$ and interpreter $g$, we define the accuracy as:
@@
 where $\left|D\right|$ denotes the cardinality of $D$, $g\left(x_{i}\right)$ is the prediction made by the interpreter for the instance $x_{i}$.
-RQ2: What predictive accuracy-related improvements does the proposed method make?
-There are two goals in validating the predictor trained by our framework (denoted as Joint_CNN): (1) The joint learning framework has had a positive impact on the multiple evaluation metrics of the predictor, and (2) its predictive performance is comparable to other notable predictors.
-For the first goal, we compare the F-measure and AUC of Joint_CNN with the DP-CNN trained not by our framework (denoted as Base_CNN).
-In addition, we incorporate three additional metrics from [58]: Percent of Perfect Cleans (PPC), Percent of Non-Perfect Cleans (PNPC), and False Omission Rate (FOR).
-These indices offer valuable insights into the framework’s ability to balance classification accuracy with practical project management benefits, such as budget savings, remaining time estimates, and failure rate predictions.
-For the second goal, we compare the F-measure and AUC of Joint_CNN with SVM, Random Forest (RF), and Deep Belief Network (DBN) based predictors.
-Additionally, although SVM, RF, and DBN are commonly used for classification tasks, they are not interpretable prediction models.
-To better demonstrate the effectiveness and advantages of our model, we also included AutoSpearman [13], a widely recognized interpretable prediction model, in the comparison.
-The Percent of Pure Cleanings (PPC) measure is defined in [58] as the percentage of correctly predicted negative cases, which indicates the amount of saved budget in the target project.
-PPC is calculated as:
-$PPC=\frac{\left|TN\right|}{\left|nt\right|}$ (14)
-where TN represents the number of true negatives and $\left|nt\right|$ denotes the total number of test instances.
-The Percent of Non-Perfect Cleans (PNPC) measure is defined in [58] as the percentage representing the sum of true positives, false positives, and false negatives over the total number of instances tested.
-This measure is defined as follows:
-$PNPC=\frac{\left|TP\right|+\left|FP\right|+\left|FN\right|}{\left|nt\right|}$ (15)
-where $\left|TP\right|$ represents the number of true positives, $\left|FP\right|$ denotes the number of false positives, and $\left|FN\right|$ indicates the number of false negatives.
-The False Omission Rate (FOR) measure is defined in [58] as the ratio of false negatives over the number of predicted cleans.
-This is used to estimate the percentage of potential failures that may occur in the target software.
-FOR is defined as:
+RQ2: What performance improvements does the joint-learning framework achieve?
+There are two goals in validating the predictor trained by the joint learning framework (denoted as Joint_CNN): (1) to assess the positive impact of the joint learning framework on predictor’s performance, and (2) to determine whether its predictive performance outperforms other notable predictors.
+For the first goal, we evaluate model performance using five key indicators: F-measure, AUC, Percent of Perfect Cleans (PPC), Percent of Non-Perfect Cleans (PNPC), and False Omission Rate (FOR).
+Although F-measure and AUC provide an overall assessment of predictive performance, they can not fully capture the effectiveness in practical SDP scenarios.
+To address this gap, we introduce PPC, PNPC, and FOR.
+For the second goal, we compare the F-measure and AUC of Joint_CNN with existing explainable and non-explainable SDP models.
+PPC, PNPC and FOR are defined as follows:
+$PPC=\frac{\left|TN\right|}{\left|D\right|}$ (14)
+$PNPC=\frac{\left|TP\right|+\left|FP\right|+\left|FN\right|}{\left|D\right|}$ (15)
 $FOR=\frac{\left|FN\right|}{\left|TN\right|+\left|FN\right|}$ (16)
-where $\left|FN\right|$ represents the number of false negatives, and $\left|TN\right|+\left|FN\right|$ is the total number of predicted cleans.
-The F-measure, also referred to as the F1-score, assesses the balance between precision and recall. It is the harmonic mean of these two values. The F-measure is defined as:
-$\color[rgb]{0,0,1}F\text{-measure}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (17)
-The Area Under the Curve (AUC) is used to evaluate the performance of binary classification models by measuring the area under its Receiver Operating Characteristic (ROC) curve, which plots the trade-off between true positive rate and false positive rate across different thresholds.
-The AUC measures the effectiveness of a classification model in distinguishing between defective and non-defective instances.
-A higher AUC indicates better discrimination, where a value closer to 1 reflects excellent predictive performance.
-RQ3: What is the global interpretability of the proposed method?
-The objectives of RQ3 are threefold: (1) to demonstrate that our approach obtains global interpretability, (2) to verify its superior computational efficiency, and (3) to demonstrate its reliability.
-Although local interpretation methods explain the decision logic for a single instance, they fail to reveal the complete structure of predictors [55].
-Global interpretable methods can fill in the gaps left by local interpretable methods.
-For the first objective, we aim to demonstrate our approach is a decision tree-like interpretable method.
-The substantial computational expense is a primary disadvantage of current global interpretation methods [56].
-For the second objective, we want to quantify the computational efficiency of our method and SP-Lime [55] by comparing the training time of them.
-The rationale for using SP-Lime as the baseline is as follows.
-First, SP-Lime is a representative example of perturbation-based global interpretable methods and most global explainable methods are based on the perturbation algorithm [55].
-Second, SP-Lime is the global version of Lime (one of baselines of RQ1).
-Selecting it as baseline is helpful to maintain the consistency of the experiment.
-RQ4: What is the effect of different features on the performance and interpretability of our approach?
-To assess the impact of individual features on the model’s performance and interpretability, we carried out an ablation experiment.
-Specifically, we took Joint_CNN as the predictor and eliminated metrics individually.
-The impacts on the predictor was evaluated by observing changes in the F-measure, AUC, and MCC, and compared these results with the model’s interpretability.
-MCC is an acronym for Matthews Correlation Coefficient that is used to measure the quality of binary classifications, particularly in imbalanced datasets.
-The MCC is defined as follows:
-$\color[rgb]{0,0,1}MCC=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (18)
-where $TP$, $TN$, $FP$ and $FN$ represent the number of true positives, true negatives, false positives, and false negatives, respectively.
-The MCC value ranges from -1 to 1, where -1 indicates complete disagreement between predictions and actual outcomes, 1 represents perfect agreement, and 0 implies performance equivalent to random guessing.
+where $\left|TP\right|,\left|FP\right|,\left|FN\right|,\left|FN\right|$ represent the number of true positives, false positives, true negative and false negatives.
+The F-measure assesses the balance between precision and recall.
+$F\text{-measure}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (17)
+The AUC is used to evaluate the performance of binary classification models by measuring the area under its Receiver Operating Characteristic (ROC) curve.
+RQ3: what’s the performance of our method in terms of global interpretability?
+The RQ3 involves three objectives: (1) evaluating the global interpretability of our approach, (2) assessing its computational efficiency, and (3) demonstrating its reliability.
+For the first objective, we aim to demonstrate that our approach provides decision tree-like interpretability.
+The decision tree is a typical global interpretation method that represents the overall decision logic through hierarchical “IF-Then” rules [45].
+Furthermore, we adopt the same quantitative method as [46] to evaluate global interpretability by measuring the contributions of metric subsets, identified by the interpreter, to the model’s predictive performance.
+Specifically, we use Joint_SDT and SP-LIME to generate six metric subsets for each dataset and use them to retrain Joint_CNN independently.
+The predictive performance of Joint_CNN with different metric subsets is measured by AUC.
+Higher AUC values indicate that the interpreter provides better global interpretability.
+For the second objective, we assess the computational efficiency of our method and SP-Lime by comparing the training times.
+Since excessive training time is a key limitation of existing global interpretation methods [47].
+For the third objective, we use three indicators, CD, FI and AI to quantify the reliability of our approach.
+RQ4: Do the metrics selected by Joint_SDT have a significant impact on the performance of Joint_CNN?
+This study evaluates the consistency between the metrics selected by Joint_SDT and those impact the performance of Joint_CNN most.
+To address this, we conduct ablation experiments in which each metric identified by Join_SDT is individually removed, and the resulting degradation in Joint_CNN’s performance (e.g., declines in Matthews Correlation Coefficient (MCC), F-measure, and AUC) is measured.
+MCC is commonly used to assess the quality of imbalanced binary classification, where the value ranges from -1 to 1.
+-1 indicates complete disagreement between predictions and ground truth, 1 represents perfect agreement.
+A strong correlation between Joint_SDT’s prioritized metrics and the magnitude of performance deterioration would further confirm that our joint-learning framework make positive contributions to both predictor and interpreter.
+$MCC=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (18)
+(a) The first interpretation
+(b) The second interpretation
+(c) The third interpretation
+(d) The fourth interpretation
+Figure 6: Multiple interpretations of LIME for a certain instance in the PC2 dataset.
 IV-A Dataset
-The 20 datasets used in the experiment were collected from the AEEEM 111 https://zenodo.org/record/3362613#.YmuGoNpByUk, NASA  [^2] http://promise.site.uottawa.ca/SERepository/datasets-page.html and PROMISE  [^3] http://promise.site.uottawa.ca/SERepository repositories.
-The effectiveness of these datasets has been demonstrated by numerous studies [45, 46, 47, 48].
+Twenty datasets used in the experiment are collected from the AEEEM 111 https://zenodo.org/record/3362613#.YmuGoNpByUk, NASA  [^2] http://promise.site.uottawa.ca/SERepository/datasets-page.html and PROMISE  [^3] http://promise.site.uottawa.ca/SERepository repositories.
+Numerous studies have demonstrated the efficacy of these datasets [48, 49, 50, 51].
 Further details about each dataset are provided in Table II.
-We observe that these datasets exhibit significant class imbalance, with defective instances accounting for only 0.41% of the PC2 dataset.
+We observe that these datasets exhibit significant class imbalance, with the highest imbalance rate reaching 0.41% from the PC2 dataset.
 TABLE II: Details of the datasets used in the experiment
@@
 IV-B Setting
-In the experiment, some parameters were set as follows: the initial learning rate $lr$ = 1e-06 and the batch size $batch\_size$ is set to 16.
-The depth of SDT $tree\_depth$ is set to 4.
-The penalty strength of SDT $penalty\_strength$ is set to 1e+1 and the penalty decay rate $penalty\_decay$ is set to 0.25.
-The hyperparameters $\gamma$, $\lambda$, $\alpha$ and $\beta$ are set to 0.8, 0.5, 1.4 and 0.6 respectively.
-The temperature hyperparameter T is set to 100.
-The window size for the Exponential Moving Average (EMA) $ema\_win\_size$ is set to 1000.
-We applied the same set of parameters across all datasets to ensure a fair comparison.
+We config our model as follows: the learning rate is initialized as 1e-06 and the batch size is set to 16.
+SGD optimizer is employed with a momentum of 0.9.
+The depth of SDT is set to 4.
+The penalty strength of SDT is set to 1e+1 and the penalty decay rate is set to 0.25.
+The hyperparameters $\alpha$ and $\beta$ in Eq.
+(2) are set to 1.4 and 0.6, and $\gamma$, $\lambda$ in Eq.
+(2) are set to 0.8 and 0.5.
+The temperature hyperparameter T in Eq.
+(6) is set to 100.
+The window size for the Exponential Moving Average (EMA) is set to 1000.
+Training stops when the model reaches the maximum number of iterations (40) or when early stopping criterias are met (no validation loss improvement for 10 consecutive epochs).
+Based on preliminary experimental results, the maximum number of iterations is set to 40, beyond which model performance saturates with minimal further improvement.
 The dataset was split into 7:2:1 for training, testing, and validation.
 The experiments are implemented by TensorFlow 2.3.0 with the NVIDIA GeForce RTX 3090 GPU.
+To ensure fairness, we applied this configuration to the experiments on all datasets.
 IV-B1 Baseline
-We utilize LIME [8] and BreakDown [28] as baselines for RQ1 and RQ3.
-We selected these techniques due to the following reasons: (1) These two methods are extensively employed by the SDP research community; (2) LIME, BreakDown, and our approach belong to feature-based interpretation techniques that strive to quantify the correlations between input features and predictive outputs; (3) The theoretical foundations of LIME and BreakDown differ substantially, thereby enriching the diversity of comparisons.
-LIME approximates the behavior of predictors in a local context, whereas BreakDown calculates the cumulative contribution of each feature to the prediction result.
+For RQ1, we utilize LIME [12] and BreakDown [30] as baselines for the following reasons: (1) These methods are extensively used in the SDP research; (2) LIME, BreakDown, and our approach belong to feature-oriented interpretation techniques that aim to quantify the correlations between input features and predictive outputs; (3) The theoretical foundations of LIME and BreakDown are different, thereby enriching the diversity of comparisons.
+For RQ2, we use SVM, Random Forest (RF), Deep Belief Network (DBN), and AutoSpearman [52] as baselines.
+These models are selected based on the following reasons: (1) SVM and RF are widely adopted non-explainable SDP models, while DBN represents DL-based approach in SDP.
+(3) AutoSpearman, in contrast, is a well-recognized explainable SDP model.
+Including it as a baseline ensures a fair comparison.
+For RQ3, we use SP-LIME [45] as the baseline for two key reasons: (1) SP-LIME and our method share the same technical foundation.
+Both of them are perturbation-based global interpretable methods.
+Moreover, SP-LIME is the representative method in this category [45].
+(2) SP-LIME is the global version of Lime (one of the baselines of RQ1).
+Selecting it as the baseline is helpful to maintain the consistency across the experiments.
 IV-B2 Data preprocessing
-We leverage normalization and different data sampling techniques to process the imbalanced data. Specifically, normalization is defined as follows:
+It involves normalization and data sampling in this section. Specifically, normalization is defined as follows:
 $\widetilde{x}=\frac{x-\min{\left(D\right)}}{\max{\left(D\right)}-\min{\left(D\right)}}$ (19)
 where $x$ is the sample of a dataset, and ${\rm min}(\cdot),{\rm max}(\cdot)$ are the minimum and maximum operations applied column-wise across the dataset.
-The features in the dataset exhibit significant differences in value range.
-Without normalization, features with larger numerical ranges might unfairly influence the model.
-Normalization ensures that all features are on the same scale, so the model treats them equally in terms of their impact.
-Moreover, normalization helps the algorithm converge faster and more smoothly.
-As shown in the Table III, normalizing the data significantly improves the model’s performance compared to using the raw features.
-We adopt two sampling techniques to the dataset: SMOTE (Synthetic Minority Oversampling Technique) and RUS (Random Undersampling).
-The technical details of SMOTE and RUS can be found in [49] and [54].
-These techniques address class imbalance in defect prediction tasks but may also affect model interpretability.
-SMOTE balances the dataset by generating synthetic samples, which may subtly alter the distribution of dataset and distort the contribution of feature to the prediction results.
-In contrast, RUS reduces the dataset size by randomly removing majority-class samples, potentially simplifying the decision process but at the risk of discarding important information, which may affect the stability of the generated explanations.
-TABLE III: Average Performance of the Models with and without Data Preprocessing Across All Datasets
-F-measure AUC
-Raw Data 0.5822 0.6041
-Preprocessed Data 0.6704 0.7109
+Normalization ensures that all features are on the same scale, and it helps the algorithm converge faster and more smoothly.
+To address the class imbalance issue, we adopt two sampling techniques: SMOTE (Synthetic Minority Oversampling Technique) [3] and RUS (Random Undersampling) [53].
 IV-B3 Statistical significance test
-To verify the significance of the improvements achieved by our approach, we introduced $Cohen^{\prime}s\ d$ in this section.
-$Cohen^{\prime}s\ d$ is used to measure the difference between two groups, which can take any value between 0 and infinity.
-When $Cohen^{\prime}s\ d$ takes values in the ranges $[0,0.2),[0.2,0.8),and\ [0.8,+\infty)$, it indicates that the differences between the two groups are small, medium, and large 444 https://en.wikipedia.org/wiki/Effect_size.
+To verify the significance of the improvements, we introduce $Cohen^{\prime}s\ d$ in this section, which takes any value between 0 and infinity.
+When $Cohen^{\prime}s\ d$ takes values in the ranges $0,0.2),[0.2,0.8),and\ [0.8,+\infty)$, it indicates that the improvement is small, medium, and large 444 https://en.wikipedia.org/wiki/Effect_size.
 The definition of $Cohen^{\prime}s\ d$ is as follows:
@@
 where $M_{1}$ and $M_{2}$ represent the means of the two group data, and $SD_{1}$ and $SD_{2}$ represent their standard deviations.
+TABLE III: The CD-10%s of LIME, BreakDown and proposed method
+LIME BreakDown Joint_SDT
+Average CD-10% 65% 74.5% 100%
+Improvement 53.85%↑ 34.23%↑ -
 TABLE IV: The AI and FI of Joint_SDT and LIME
-$AI$ $FI$
+AIAI FIFI
 LIME Joint_SDT LIME Joint_SDT
@@
 $Cohen^{\prime}s\ d$ 0.6792 - 0.9451 -
-TABLE V: Average CDs of LIME, BreakDown and proposed method
-LIME BreakDown Joint_SDT
-Average 65% 74.5% 100%
-Improvement 53.85%↑ 34.23%↑ -
-IV-C Experimental results
-The experimental results are presented as follows.
-IV-C1 The results of RQ1
-To evaluate the consistency of the interpreter trained by our framework (denoted as Joint_SDT), LIME and BreakDown, we employed each of them to interpret an instance ten times, and calculate the average CD for each method.
-As shown in Table V, the average CD-10% of LIME, BreakDown and Joint_SDT are 65%, 74.5%, and 100%.
-This result suggests that LIME exhibits poor consistency.
-To further investigate the consistency of LIME, we visualized its outputs in Fig.
-6 and calculated the coefficient of variation (CV) for the weights assigned to different metrics across the ten interpretations.
-The results show that the CV is 42.18%.
-The high CV suggests that LIME is unable to consistently assign similar weights to the same metric across multiple explanations.
-We believe this inconsistency is partly due to the sampling technique and the hyperparameters used by LIME.
-(a) The first interpretation
-(b) The second interpretation
-(c) The third interpretation
-(d) The fourth interpretation
-(e) The fifth interpretation
-(f) The sixth interpretation
-(g) The seventh interpretation
-(h) The eighth interpretation
-(i) The ninth interpretation
-(j) The tenth interpretation
-Figure 6: Multiple interpretations of LIME for a certain instance in the PC2 dataset.
-Furthermore, we conducted a comparison between the interpretations generated by LIME and Joint_SDT in terms of AI and FI.
-Since BreakDown does not provide an explicit surrogate function like LIME [8], its AI and FI of BreakDown cannot be calculated.
-As shown in Table IV, across all projects, Joint_SDT achieves an average improvement of 12.38% in AI and 19.45% in FI compared to LIME.
-The $Cohen^{\prime}s\ d$ results in the table further confirm that Joint_SDT outperforms LIME, with a large effect size in AI and a moderate effect size in FI.
-IV-C2 The results of RQ2
-TABLE VI: Prediction Performance of Joint_CNN and Base_CNN in Terms of F-measure and AUC
-F-measure AUC
-Base_CNN Joint_CNN Base_CNN Joint_CNN
-JDT 0.6802 0.8544 0.7297 0.8360
-LC 0.5686 0.8285 0.6923 0.6862
-ML 0.5686 0.7123 0.6897 0.6418
-PDE 0.7475 0.7111 0.7039 0.7032
-jm1 0.5755 0.8621 0.6240 0.6280
-kc1 0.6443 0.7303 0.7545 0.8087
-mc1 0.7352 0.8621 0.8082 0.6280
-PC1 0.7475 0.8621 0.7039 0.6280
-PC2 0.4996 0.8621 0.5000 0.6280
-PC3 0.5588 0.7957 0.7634 0.8324
-PC4 0.7021 0.7472 0.8169 0.7843
-ant-1.7 0.7991 0.7907 0.8276 0.8022
-camel-1.6 0.6932 0.7337 0.5992 0.6687
-ivy-1.2 0.8273 0.8347 0.6016 0.8594
-jedit-4.1 0.5862 0.6946 0.5429 0.8125
-log4j-1.0 0.7755 0.6520 0.8500 0.8500
-lucene-2.4 0.5623 0.6947 0.5482 0.7643
-poi-3.0 0.7156 0.6951 0.7263 0.7414
-synapse1.2 0.6616 0.6486 0.8235 0.7516
-xerces-1.3 0.7921 0.8407 0.8498 0.7875
-Average 0.6720 0.7706 0.7078 0.7421
-Improvement 14.67%↑ - 4.85%↑ -
-$Cohen^{\prime}s\ d$ 1.1321 - 0.3553 -
-TABLE VII: Average Performances of Joint_CNN and Base_CNN in Terms of PPC, PNPC and FOR
-Group PPC PNPC FOR
-Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN
-AEEEM 0.6570 0.7394 0.3430 0.2606 0.2597 0.1662
-NASA 0.6811 0.7708 0.3189 0.2292 0.0503 0.0256
-PROMISE 0.4974 0.5167 0.5026 0.4833 0.1867 0.1398
-Average 0.6118 0.6756 0.3882 0.3244 0.1656 0.1105
-Improvement 10.43% - 19.67$\downarrow$ - 49.86$\downarrow$ -
-According to Table VI, we observed that Joint_CNN outperforms Base_CNN in terms of both F-measure and AUC.
-Across all datasets, Joint_CNN achieves an average improvement of 14.67% in F-measure and 4.85% in AUC compared to Base_CNN.
-Table VII presents the average performance of Base_CNN and Joint_CNN on three groups: PROMISE, NASA, and AEEEM, evaluated using the metrics PPC, PNPC, and FOR.
-The results in Table VII show that, compared to Base_CNN, Joint_CNN achieves improvements of 10.43% in PPC, 19.67% in PNPC, and 49.86% in FOR.
-This indicates that Joint_CNN is more accurate in predicting clean modules, more effective in predicting defective instances on average, and better at reducing false negative instances across the target projects.
-The $Cohen^{\prime}s\ d$ results indicate a large effect size in F-measure and a moderate effect size in AUC, further confirming the superiority of Joint_CNN over Base_CNN.
-According to Table VIII, joint_CNN outperforms SVM, RF, DBN and AutoSpearman in terms of both F-measure and AUC.
-Specifically, Joint_CNN achieves average improvements of 42.86%, 30.63%, 61.75% and 2.13% in F-measure compared to SVM, RF, DBN and AutoSpearman, respectively.
-Similarly, in AUC, Joint_CNN demonstrates average improvements of 10.93%, 15%, 20.75% and 0.97%.
-The $Cohen^{\prime}s\ d$ results further validate the performance advantages of Joint_CNN.
-In terms of F-measure, the effect sizes are large, indicating significant improvements over SVM, RF and DBN.
-While the effect size for AutoSpearman is smaller, Joint_CNN still demonstrates overall better performance.
-For AUC, the effect sizes vary across different comparisons.
-Specifically, the effect size for SVM is medium, while the effect sizes for RF and DBN are large, indicating significant improvements.
-Although the effect size for AutoSpearman is small, Joint_CNN still demonstrates overall robustness.
-TABLE VIII: The F-measure and AUC performance of different models on all datasets
+TABLE V: Prediction Performance of Joint_CNN and Base_CNN in Terms of F-measure, AUC, PPC, PNPC and FOR
+F-measure AUC PPC PNPC FOR
+Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN
+JDT 0.6802 0.8544 0.7297 0.8360 0.6000 0.6900 0.4000 0.3100 0.2683 0.1585
+LC 0.5686 0.8285 0.6923 0.6862 0.7714 0.7857 0.2286 0.2143 0.1692 0.1538
+ML 0.5686 0.7123 0.6897 0.6418 0.6203 0.6952 0.3797 0.3048 0.3216 0.2398
+PDE 0.7475 0.7111 0.7039 0.7032 0.7000 0.7867 0.3000 0.2133 0.2105 0.1128
+jm1 0.5755 0.8621 0.6240 0.6280 0.5417 0.8036 0.4583 0.1964 0.1335 0.0012
+kc1 0.6443 0.7303 0.7545 0.8087 0.5024 0.5308 0.4976 0.4692 0.0783 0.0345
+mc1 0.7352 0.8621 0.8082 0.6280 0.8786 0.7818 0.1214 0.2182 0.0012 0.0652
+PC1 0.7475 0.8621 0.7039 0.6280 0.6909 0.8182 0.3091 0.1818 0.0500 0.0816
+PC2 0.4996 0.8621 0.5000 0.6280 0.9911 0.9875 0.0089 0.0125 0.0018 0.0018
+PC3 0.5588 0.7957 0.7634 0.8324 0.7707 0.6752 0.2293 0.3248 0.0397 0.0093
+PC4 0.7021 0.7472 0.8169 0.7843 0.8082 0.6644 0.1918 0.3356 0.0167 0.0300
+ant-1.7 0.7991 0.7907 0.8276 0.8022 0.7733 0.5733 0.2267 0.4267 0.2267 0.0444
+camel-1.6 0.6932 0.7337 0.5992 0.6687 0.5567 0.5876 0.4433 0.4124 0.1692 0.1739
+ivy-1.2 0.8273 0.8347 0.6016 0.8594 0.7778 0.7500 0.2222 0.2500 0.0667 0.1000
+jedit-4.1 0.5862 0.6946 0.5429 0.8125 0.5312 0.5938 0.4688 0.4062 0.0556 0.0500
+log4j-1.0 0.7755 0.6520 0.8500 0.8500 0.5714 0.3571 0.4286 0.6429 0.2727 0.0235
+lucene-2.4 0.5623 0.6947 0.5482 0.7643 0.2941 0.3529 0.7059 0.6471 0.4737 0.3684
+poi-3.0 0.7156 0.6951 0.7263 0.7414 0.4000 0.5333 0.6000 0.4667 0.3793 0.1724
+synapse1.2 0.6616 0.6486 0.8235 0.7516 0.3077 0.6154 0.6923 0.3846 0.2000 0.1111
+xerces-1.3 0.7921 0.8407 0.8498 0.7875 0.6304 0.6957 0.3696 0.3043 0.1212 0.1111
+Average 0.6720 0.7706 0.7078 0.7421 0.6359 0.6639 0.3641 0.3361 0.1628 0.1022
+Improvement 14.67%↑ - 4.85%↑ - 4.4%↑ - 8.33% ↑ - 59.30%↑ -
+$Cohen^{\prime}s\ d$ 1.1321 - 0.3553 - 0.0763 - 0.1665 - 0.5338 -
+TABLE VI: The F-measure and AUC performance of different models on all datasets
 F-measure AUC
@@
 $Cohen^{\prime}s\ d$ 2.3697 1.7196 2.9703 0.1712 - 0.7337 0.9543 1.1906 0.0657 -
-We believe there are two reasons related to the improvement: (1) We explicitly integrate the prediction error $L_{pred}(f,D)$ into the loss function of the joint learning, enabling the predictor to capture discriminative information contained in the data during the training process, and subsequently improving the corresponding accuracy.
-(2) The $L_{int}$ can be regarded as a penalty term of $L$, which effectively encourages the framework to learn more generalizable representations of the data, leading to improved predictive accuracy for the data that is not contained in training dataset.
+TABLE VII: The CD, FI and AI values of Joint_SDT and SP-Lime
+Method CD FI AI
+Joint_SDT 1.00 0.8714 0.8429
+SP-Lime 0.81 0.7933 0.8012
+IV-C Experimental results
+The experimental results are presented as follows.
+IV-C1 The results of RQ1
+To evaluate the consistency of Joint_SDT, LIME and BreakDown, we applied each of them to interpret the same instance ten times.
+As shown in Table III, the average CD-10% scores of LIME, BreakDown and Joint_SDT are 65%, 74.5%, and 100%.
+This result suggests that both LIME and BreakDown exhibit poor consistency.
+To further investigate the root cause of such poor consistency, we visualized the interpretations of LIME in Fig.
+6.
+We can see that LIME is unable to assign consistent weights to the same metric across multiple explanations, which we believe it is the root of poor consistency.
+Moreover, we conducted a comparison between the interpretations generated by LIME and Joint_SDT in terms of AI and FI.
+Since BreakDown does not provide an explicit surrogate function like LIME [12], we are unable to calculate its AI and FI.
+As shown in Table IV, Joint_SDT achieves an average improvement of 12.38% in AI and 19.45% in FI compared to LIME.
+Furthermore, the $Cohen^{\prime}s\ d$ results indicate that the significance of improvement in AI is medium, while in FI, it is large.
+This results further confirm that Joint_SDT outperforms LIME in the consistency of interpretation.
+Although Joint_SDT generally shows improvements in AI and FI, its performance is worse than LIME on certain datasets, such as Log4j.
+This is largely due to the limited amount of data, where sparse distributions enable LIME to maintain consistent interpretability by reducing sampling randomness.
+However, the neural network-like structure of Joint_SDT requires a larger volume of data for stable and reliable feature importance estimation.
+As a result, Joint_SDT struggles to achieve desirable performance on small datasets.
+IV-C2 The results of RQ2
+According to Table V, we observe that Joint_CNN outperforms Base_CNN in terms of F-measure, AUC, PPC, PNPC, and FOR on most datasets, and the average improvements are 14.67%, 4.85%, 4.4%, 8.33% and 59.30% respectively.
+The corresponding $Cohen^{\prime}s\ d$ test results indicate that the effect sizes of the improvements in the five indicators are large, medium, small, small, and medium.
+From these results, we can draw the following conclusions: (1) The improvement of Joint_CNN over Base_CNN across all evaluation metrics indicates that the joint-learning framework makes a positive contribution in improving the performance of SDP models.
+(2) The improvements in PNPC and FOR suggest that Joint_CNN is more reliable in distinguishing defective and non-defective instances, which is crucial for real-world SDP applications.
+(3) The constant improvements across all datasets suggests that the joint-learning framework generalize well and is not limited to specific dataset.
+Furthermore, according to Table VI, joint_CNN outperforms SVM, RF, DBN and AutoSpearman in both F-measure and AUC.
+Specifically, Joint_CNN achieves average improvements of 42.86%, 30.63%, 61.75% and 2.13% in F-measure and 10.93%, 15%, 20.75% and 0.97% in AUC.
+The $Cohen^{\prime}s\ d$ test results indicate that the effect sizes of of improvements in F-measure are large, large, large, and small, while those in AUC are medium, large, large, and small.
+According to these results, we can confirm that Joint_CNN achieves better accuracy and demonstrates strong robustness across all dataset.
+We believe there are two reasons related to the improvements: (1) We explicitly integrate the prediction error $L_{pred}(f,D)$ into the loss function of the joint-learning framework, guiding the predictor to capture discriminative information contained in the data.
+(2) The revised distillation loss $L_{int}$ can be regarded as a penalty term of $L$, which effectively encourages the framework to learn more generalized features of the data, improving the model’s robustness.
 IV-C3 The results of RQ3
-To demonstrate the global interpretability of our approach, we take the PC2 dataset as an example.
+To demonstrate the global interpretability of our approach, we use the PC1 dataset as an example.
 As presented in Fig.
 5, the decision logic can be abstracted as the path from the “DECISION_DENSITY” metric (denoted as the root) to the leaf node (denoted as 1 or 0).
-For example, one feasible decision path can be presented as DECISION_DENSITY -> MAINTENANCE_SEVERITY -> PERCENT_COMMENTS -> PERCENT_COMMENTS -> 0, which is similar to the form of decision tree’s interpretation.
-So, we can conclude our approach achieves the global interpretability.
+For example, one feasible decision path can be presented as DECISION_DENSITY -> MAINTENANCE_SEVERITY -> PERCENT_COMMENTS -> PERCENT_COMMENTS -> 0, which is similar to the interpretation format of a decision tree.
+Furthermore, following the framework of [46], we visualized the AUCs of Joint_CNNs using different metric subsets identified by Joint_SDT and SP-LIME in Fig.
+7.
+The results demonstrate that our method achieves higher AUC values.
+This confirms that the metric subsets selected by Joint_SDT are more effective for explaining the model’s global decision-making logic compared to those selected by SP-LIME.
 We measure the computational cost by calculating the average training time.
@@
 Furthermore, we calculate the CD, FI and AI values for LC dataset to evaluate the reliability of our approach and SP-Lime.
-As presented in table IX, our method achieves a better performance in all three indicators, indicating our method is more reliable than SP-Lime.
+As presented in table VII, our method achieves a better performance in all three indicators.
 The reason for this result can be partially attributed to the random sampling involved in perturbation processes of SP-Lime [45].
-TABLE IX: The CD, FI and AI values of Joint_SDT and SP-Lime
-Method CD FI AI
-Joint_SDT 1.00 0.8714 0.8429
-SP-Lime 0.81 0.7933 0.8012
-IV-C4 The results of RQ4
-We conducted an ablation experiment on 5,589 instances in the PC2 dataset to analyze the impact of different features on model performance.
-The interpretations generated by our approach are presented in Fig.
-5.
-In the interpretations of our approach, the metrics that have a greater impact on the prediction results are located closer to the root node.
-Meanwhile, the impacts on the F-measure, AUC, and MCC after removing certain metrics are presented in Fig.
-7.
-We can see that the accuracy of Joint_CNN went down significantly across MCC, F-measure and AUC three indicators after removing the “DECISION_DENSITY”.
-This indicates that this metric is critical for the accuracy of the predictor.
-This finding is consistent with the interpretation of our approach, as our approach placed it at the root node of SDT, indicating its decisive role in the model’s decision-making process.
-Thus, the significant decline in performance after removing this feature further corroborates its key role in the model.
-The importance of this feature is strongly evidenced both by the performance change and its position in the explanation.
-To enhance the comprehensiveness of the analysis, the ablation experiment was extended to all datasets.
-The most important features interpreted by our approach for each dataset, as well as the ranking of performance degradation on the F-measure, AUC, and MCC after removing these features, are presented in Table X.
-The results demonstrate that the performance degradation is consistent with our explanatory results, further validating that our method effectively captures key features to improve prediction performance and provide accurate model interpretations.
-TABLE X: Ranking of Feature Importance Based on SDT Interpretations and Ablation Results
-Dataset Top Feature Identified by SDT F-measure Rank AUC Rank MCC Rank
-JDT WCHU_numberOfMethodsInherited 4 9 5
-LC WCHU_cbo 2 1 1
-ML ck_oo_numberOfPublicMethods 5 9 5
+Figure 7: Comparison of AUC values of Joint_CNNs using metric subsets selected by Joint_SDT and SP-LIME on LC, PC1, and Lucene datasets.
+(a) F-measure
+(b) AUC
+(c) MCC
+Figure 8: Ablation study results across multiple indicators. The gray dashed line indicates the performance of DP-CNN before removing metrics.
+TABLE VIII: Rankings of Feature Importance
+Dataset The Most Important Feature Identified by Joint_SDT F-measure Rank AUC Rank MCC Rank
+JDT WCHU_numberOfMethodsInherited 4 3 5
+LC WCHU_cbo 2 1 2
+ML ck_oo_numberOfPublicMethods 3 2 5
 PDE WCHU_fanOut 3 3 5
-jm1 HALSTEAD_DIFFICULTY 7 7 1
+jm1 HALSTEAD_DIFFICULTY 3 2 1
 kc1 v 1 2 2
 mc1 LOC_TOTAL 1 1 1
-PC1 branchCount 2 1 1
+PC1 branchCount IV 2 1 1
 PC2 DECISION_DENSITY 1 4 1
@@
 camel-1.6 lcom 4 3 4
-ivy-1.2 cam 2 7 4
-jedit-4.1 lcom3 6 10 4
+ivy-1.2 cam 2 5 4
+jedit-4.1 lcom3 4 5 4
 log4j-1.0 max_cc 4 4 3
@@
 synapse-1.2 cbo 1 1 4
-xerces-1.3 max_cc 2 8 2
-(a) F-measure
-(b) AUC
-(c) MCC
-Figure 7: Ablation study results across multiple indicators. The gray dashed line indicates the performance of DP-CNN before removing metrics.
+xerces-1.3 max_cc 2 5 2
+IV-C4 The results of RQ4
+According to the structure of Joint_SDT, the most important metric is positioned at the root of the SDT.
+For instance, in the PC2 dataset, ”DECISION_DENSITY” is identified as the most important metric.
+To validate this finding, we conduct an ablation experiment that removes each metric selected by Joint_SDT individually and the corresponding changes are observed in the performance of Joint_CNN.
+The ranking of performance degradation reflects the importance of each metric to the model.
+The ablation experiment results for PC2 are visualized in Fig.
+8.
+We can see that removing ”DECISION_DENSITY” results in the most significant performance decline across the three indicators: MCC, F-measure, and AUC.
+This confirms that ”DECISION_DENSITY” is the most important metric for PC2 dataset.
+This aligns with the interpretation of our method.
+To enhance the comprehensiveness of the analysis, we extended the ablation experiment to all datasets.
+Table VIII presents the most important metrics identified by Joint_SDT across the twenty dataset.
+The results show that the root metrics identified by Joint_SDT typically correspond to the most significant performance degradations in the ablation experiments.
 V Threats to validity
-While the proposed model demonstrates advancements in both prediction accuracy and interpretability, certain limitations may affect its generalizability and reliability. The details are as follows.
-  • Choice of dataset: The experiment was conducted on the publicly available datasets.
-These datasets may not fully capture the diversity of software systems in real-world scenarios, potentially limiting the robustness of the model.
-  • Hyperparameter setting: The model’s performance is related with the selection of hyperparameters for both the predictor and interpreter.
-Suboptimal tuning may lead to overfitting or underfitting, resulting in less accurate predictions or interpretations.
-  • Application scenarios: In this paper, we focus on the within-project defect prediction that the training data and the test data are generated from the same project.
-However, the performance of our model in the context of the cross-project setting is not evaluated.
+While the proposed model demonstrates improvements in both prediction accuracy and interpretability, certain limitations may affect its generalizability. The details are as follows.
+  • Choice of dataset: The experiments were conducted on the publicly available datasets.
+These datasets may not fully capture the diversity of software systems in real-world scenarios, potentially limiting the generalizability of the model.
+  • Hyperparameter setting: The model’s performance is related with the setting of hyperparameters for both the predictor and interpreter.
+Suboptimal setting may lead to less accurate predictions or interpretations.
+  • Application scenarios: In this paper, we focus on the within-project SDP. However, the performance of our model in the context of the cross-project setting has not been evaluated.
 VI Conclusion and future works
 In this paper, we explore the possibility of designing a defect predictor and its corresponding interpreter collaboratively.
-Unlike most existing approaches, which treat defect prediction and interpretation as separate tasks, we consider them strongly correlated.
-Based on this assumption, we introduce a framework aimed at improving the reliability of interpretation while simultaneously enhancing predictive accuracy by utilizing interpretation results.
-Based on the proposed framework, we conducted extensive empirical evaluations, which revealed the following key findings:
-  • Compared to existing interpretation methods such as LIME and BreakDown, the proposed framework significantly improves the consistency and reliability of interpretations, providing robust explanations from a local perspective.
-This improvement is attributed to the incorporation of the Knowledge Distillation (KD) principle, which enables the accurate transfer of decision-making knowledge from the predictor to the interpreter.
-  • The framework demonstrates superior performance on widely used datasets in terms of both F-measure and AUC, outperforming baseline models such as SVM, RF, DBN, and Base_CNN.
-By explicitly incorporating interpretation results into the loss function, the framework captures more discriminative information, leading to more robust predictive outcomes.
-  • The proposed method offers global interpretability, providing developers with transparent global insights and decision-making rationales.
-This capability ensures that the framework not only achieves high accuracy but also delivers comprehensive explanatory support for real-world defect prediction tasks.
-  • Ablation analysis further demonstrates that the framework effectively identifies key features that have the greatest impact on prediction outcomes, showcasing its ability to provide reliable explanations that align closely with the actual importance of these features.
-Future research will investigate several possibilities to further improve the proposed model, including the following:
-  • Expand the research by including a wider array of datasets that accurately reflect real-world software systems. Evaluate the robustness and generalizability of our model accordingly.
-  • Incorporate automatic hyperparameter optimization methods into our model to reduce the likelihood of suboptimal tuning.
-  • Assess the model’s performance in several defect prediction scenarios, such as the cross-project defect prediction.
+Unlike most existing approaches, which treat defect prediction and interpretation as separate tasks, we regard them as strongly correlated.
+We introduce a framework aiming at improving the reliability of interpretation and enhancing predictive accuracy simultaneously.
+Based on the extensive empirical evaluations, we obtained the following key findings:
+  • Compared to existing interpretation methods such as LIME and BreakDown, the proposed framework significantly improves the reliability of interpretations through the incorporation of the KD principle.
+  • The joint-learning framework demonstrates superior accuracy on widely used datasets by explicitly incorporating interpretation results into the loss function.
+Future research will explore the following directions:
+  • Using diverse real-world datasets to evaluate the robustness and generalization of our method.
+  • Integrating automatic hyperparameter optimization to reduce manual tuning efforts.
+  • Assess the model’s performance in different defect prediction scenarios, such as cross-project defect prediction.